### v2

# Asymptotic behavior of eigenvalues of large rank perturbations of large random matrices

Abstract
The paper is concerned with deformed Wigner random matrices. These matrices are closely connected with Deep Neural Networks (DNNs): weight matrices of trained DNNs could be represented in the form $R+S$, where $R$ is random and $S$ is highly correlated. The spectrum of such matrices plays a key role in rigorous underpinning of the novel pruning technique based on Random Matrix Theory. Mathematics has been done only for finite-rank matrix $S$. However, in practice rank may grow. In this paper we develop asymptotic analysis for the case of growing rank.
1 Introduction
Random Matrix Theory (RMT) is a classical theory that has been developing for more than 70 years. Initially, RMT arose from problems in nuclear physics and found its applications in mathematics, physics, finance, and many other disciplines. Recently, new problems have been arising from the area of Machine Learning. Indeed, often the weight matrices of Deep Neural Networks (DNNs) are initialized randomly. Moreover, modern DNNs have large weight matrices, which is why their spectral properties can be described by asymptotic behavior of $N\times N$ random matrices as $N$ goes to infinity.
1.1 Formulation of the problem
Martin and Mahoney in [4] showed that weight matrices of trained DNNs can be modeled by large $N\times M$ random matrices of the form
$W=\frac{1}{\sqrt{N}}R+S,$ (1.1)
where $R$ is a rectangular matrix whose entries are i.i.d. random variables with zero mean and variance $\sigma^{2}$ and $S$ is a real symmetric non-random (or random, but strongly correlated) matrix. The present paper is concerned with similar but simpler $N\times N$ deformed Gaussian Orthogonal Ensemble matrices of the same form
$W=\frac{1}{\sqrt{N}}R+S,$ (1.2)
where $R$ is restricted to be a quadratic real symmetric matrix whose entries are Gaussian random variables $\mathcal{N}(0,\sigma^{2})$ and $S$ is also a real symmetric non-random matrix.
The empirical spectral distribution (ESD, a.k.a. the Normalized Counting Measure) for an $N\times N$ symmetric matrix $X$ is defined as:
$\mu_{N}(\Delta)=\frac{1}{N}\sum_{j=1}^{N}\mathbbm{1}_{\Delta}(\lambda_{j})$ (1.3)
where $\Delta$ is an interval in $\mathbb{R}$, $\lambda_{j}$ – eigenvalues of $X$ and $\mathbbm{1}_{\Delta}$ is a characteristic function of the set $\Delta$. For a probability measure $\tau$ on $\mathbb{R}$, denote by $g_{\tau}$ its Stieltjes transform defined for $z\in\mathbb{C}\setminus\operatorname{supp}(\tau)$ by
$g_{\tau}(z)=\int\limits_{\mathbb{R}}\frac{d\tau(t)}{t-z}.$ (1.4)
In 1972 Pastur established that the ESD $\mu$ of the matrices (1.2) converges as $N\to\infty$, assuming the ESD $\nu$ of $S$ converges to the measure $\nu_{0}$ [5], see also [1]. It was also shown that the Stieltjes transform $g_{\mu_{0}}(z)$ of the limiting ESD $\mu_{0}$ of the matrix $W$ defined (1.2) satisfies the equation
$g_{\mu_{0}}(z)=g_{\nu_{0}}(\omega_{\mu_{0}}(z)),\quad z\in\mathbb{C}_{+},$ (1.5)
where for any measure $\tau$ we denote
$\omega_{\tau}(z)=z+\sigma^{2}g_{\tau}(z).$ (1.6)
In the particular case when $\nu_{0}$ is a delta-measure at zero, $\mu_{0}$ coincides with a well-known semicircular law $\mu_{sc}$ [8]
$\frac{d\mu_{sc}}{dx}(x)=\frac{1}{2\pi\sigma^{2}}\sqrt{4\sigma^{2}-x^{2}}\mathbbm{1}_{[-2\sigma,2\sigma]}(x).$ (1.7)
Moreover, if rank of $S$ is finite, i.e. does not depend on $N$, Péché established the behavior of the largest eigenvalue $\lambda_{1}(W)$ in [7]. It was shown that if $\lim\limits_{N\to\infty}\lambda_{1}(S)=\theta_{1}$ then almost surely
$\lim\limits_{N\to\infty}\lambda_{1}(W)=\begin{cases}2\sigma,\quad\text{if }\theta_{1}<\sigma,\\ \Phi(\theta_{1})=\theta_{1}+\frac{\sigma^{2}}{\theta_{1}},\quad\text{if }\theta_{1}\geq\sigma,\end{cases}$ (1.8)
where
$\Phi(z)=z-\sigma^{2}g_{\nu_{0}}(z).$ (1.9)
Let $\mu$ and $\nu$ be the ESDs of $W$ and $S$ respectively. We impose the following assumptions on the matrix $S$ and its spectrum.
Assumption 1.
As $N\to\infty$, the ESD $\nu$ weakly converges to the measure $\nu_{0}$.
Assumption 2.
There are $r=r(N)$ eigenvalues of $S$ outside $\operatorname{supp}\nu_{0}$. Let $\lim\limits_{N\to\infty}r(N)=\infty$ and $r(N)=o(N)$ as $N\to\infty$.
Assumption 3.
As $N\to\infty$, the measure $\frac{N}{r}(\nu-\nu_{0})$ weakly converges to the signed measure $\nu_{1}$, $\nu_{1}(\mathbb{R}\setminus\operatorname{supp}\nu_{0})=-\nu_{1}(\operatorname{supp}\nu_{0})=1$. Moreover, $\lim\limits_{N\to\infty}\frac{N}{r}\int t^{2}d(\nu-\nu_{0})(t)=\int t^{2}d\nu_{1}(t)$.
The crucial distinction from the previous results is Assumption 2. All the previously obtained theorems assume that the deterministic matrix $S$ has a finite rank $r(N)=r^{\prime}$ independent of $N$.
In [3] Berlyand et al provided mathematical underpinning to the MP-pruning algorithm in the following setting: the weight matrix to be pruned can be represented as a sum of random and low-rank (‘‘signal’’) components with non-intersecting spectra (1.1). The asymptotical behavior of the spectrum of the matrices (1.1) plays a key role in bounding the difference between weight matrices before and after pruning, although the mathematics is done only for finite rank. However, restriction $\lim_{N\to\infty}r(N)<\infty$ severe bounds applicability of pruning. In practice it seems that a growing rank $r(N)\to\infty$, as $N\to\infty$, is in better agreement with numerically computed weight matrices of DNNs. Numerical simulations show that the number of singular values of weight matrices may grow, see Subsection 1.2 and Figure 1. This paper develops asymptotical analysis for the case of growing rank.
1.2 Motivation: weight matrices of DNNs
DNNs is a broad class of computational algorithms inspired by the human brain. They are used for classification, approximation, and image recognition, among others. DNNs have become central to modern machine learning, and understanding their mathematical properties remains an important challenge.
Let us consider a DNN for solving a classification problem. The setup of a classification problem is as following. There is a set of some objects $\mathcal{S}$. Each object $s\in\mathcal{S}$ belongs to one of $K$ classes. Given the object, one wants to determine its class. We look for a solution in the form of a function $\varphi\colon\mathcal{S}\to[0;1]^{K}$, interpreting components of $\varphi(s)$ as probabilities for $s$ to belong to each class.
From the mathematical point of view, a feedforward fully-connected DNN for solving the classification problem is a function of the following form:
$\varphi(\cdot,\alpha)=\sigma\circ\lambda\circ f_{M}(\cdot,\alpha^{M})\circ\lambda\circ f_{M-1}(\cdot,\alpha^{M-1})\circ\cdots\circ\lambda\circ f_{1}(\cdot,\alpha^{1})$ (1.10)
where each $f_{k}:\mathbb{R}^{N_{k-1}}\to\mathbb{R}^{N_{k}}$, $f_{k}(x,\alpha^{k})=W_{k}x+\beta_{k}$ is an affine function, $W_{k}$ is a $N_{k}\times N_{k-1}$ weight matrix, $\beta_{k}\in\mathbb{R}^{N_{k}}$ is a bias vector, $\lambda:\mathbb{R}\to\mathbb{R}$ is a nonlinear activation function, which is applied component-wise (for example, absolute value), $\sigma$ is a function which normalizes vector’s components to probabilities. The entries of $W_{k}$ and the biases $\beta_{k}$ are the parameters $\alpha^{k}$ of a DNN. The parameters are optimized by minimizing the loss function, using, for example, stochastic gradient descent. See [2] for more details.
Pruning parameters of a DNN is the process of removing certain weights (set them equal to zero) from the network in a way which preserves the overall performance. The goal is to eliminate components that contribute little to the output of the DNN function, resulting in a more sparse and more efficient architecture. Pruning parameters of deep neural networks is an important task because it reduces complexity of the function, leading to smaller, faster, and more efficient models. This not only decreases memory and computational requirements — crucial for deployment on resource-constrained devices — but can also improve generalization by preventing overfitting. In [3] Berlyand et al provided mathematical underpinning to the MP-pruning algorithm in the following setting: the weight matrix to be pruned can be represented as a sum of random and low-rank (‘‘signal’’) components with non-intersecting spectra (1.1). The asymptotical behavior of the spectrum of the matrices (1.1) plays a key role, although the mathematics is done only for finite rank.
The fixed low-rank structure of the ‘‘signal’’ matrix allows for certain analytical tools to be employed, leading to provable guarantees on network behavior, particularly the bounds on accuracy and loss after pruning. However, numerical experiments reveal that this low-rank assumption does not always hold in practice. Realistic weight matrices often have a spectrum that has a ‘‘signal’’ part with the number of non-zero eigenvalues increasing with respect to the matrix size (see Figure 1). See Appendix A for more details. These observations suggest that the existing theoretical results, while insightful, are not fully representative of the behavior seen in real-world networks.
Figure 1: Numeric simulation of a DNN with 3 layers. Show the dependence of the number of non-zero eigenvalues of the ‘‘signal’’ matrix, so called spikes, on the size of the matrix.
In this work, we study a mathematical problem applicable in a simplified yet insightful setting: a DNN with two consecutive layers of equal size. This structure implies that the weight matrix connecting these layers is square. Furthermore, we impose an additional assumption of symmetry on the weight matrix, which not only aligns with certain model architectures but also simplifies the theoretical analysis.
Motivated by the gap between existing rigorous proofs and numerical results, our goal is to generalize the existing theoretical framework to accommodate symmetric weight matrices without a fixed-rank constraint. By relaxing the low-rank assumption, we aim to provide a more comprehensive understanding of the network’s accuracy and the behavior of the loss function. This generalization not only brings the theory closer to practical scenarios but also raises interesting mathematical questions about the behavior of random matrices beyond the low-rank-perturbation setting. Through this work, we seek to bridge the divide between theory and numerical computations, contributing to a more robust foundation for analyzing deep learning algorithms.
2 Main results
In this section we present our main results. We first consider the measure $\frac{N}{r}(\mu-\mu_{0})$, which may be interpreted as an ESD of eigenvalues outside the bulk. The following theorem proves existence of the limiting measure and describes it.
Theorem 2.1.
Let $W$ be a random matrix of the form (1.2). Let $\mu$ and $\nu$ be the ESDs of matrices $W$ and $S$, respectively, $\mu_{0}$ be the limit of $\mu$ as $N\to\infty$, and let Assumptions 1–3 hold. Then a signed measure $\frac{N}{r}(\mu-\mu_{0})$ converges weakly in distribution, as $N\to\infty$, to a non-random measure $\mu_{1}$. Moreover, if $\mu_{1}$ has a density, then for any measurable set $\Delta$ which does not intersect with $\operatorname{supp}\mu_{0}$
$\mu_{1}(\Delta)=\nu_{1}(\omega_{\mu_{0}}(\Delta)).$ (2.1)
The next theorem describes the asymptotic behavior of individual eigenvalues. To introduce the theorem, we need more notation. For any matrix $X$ let its eigenvalues be
$\lambda_{1}(X)\geq\lambda_{2}(X)\geq\dotsb\geq\lambda_{N}(X).$ (2.2)
Also, denote the positive part of $\nu_{1}$ by $\nu_{1}^{+}$.
Theorem 2.2.
Let $W$ be a random matrix of the form (1.2). Let $\mu$ and $\nu$ be the ESDs of matrices $W$ and $S$ respectively and let Assumptions 1–3 hold. Additionally, we assume $\operatorname{supp}\nu_{1}^{+}=[a,b]$, $\operatorname{dist}(a,\operatorname{supp}\nu_{0})>0$, $\lambda_{j}(S)$, $1\leq j\leq r$, are outside $\operatorname{supp}\nu_{0}$ and
$\lim_{N\to\infty}\max_{1\leq j\leq N}\operatorname{dist}(\lambda_{j}(S),\operatorname{supp}\nu_{0}\cap[a,b])=0;$ (2.3)
$\lim_{N\to\infty}\max_{1\leq j\leq r}\operatorname{dist}(\lambda_{j}(S),[a,b])=0.$ (2.4)
Then for any $j(N)\leq r(N)$
$\lambda_{j(N)}(W)-\Phi(\lambda_{j(N)}(S))\to 0,\text{ as }N\to\infty$ (2.5)
in probability, where $\Phi(x)$ is defined in (1.9).
Remarks
  1. 1.
[…]
  2. 2.
    Theorem 2.2 agrees well with the result (1.8) obtained by Péché [7].
The paper is organized as follows. Section 3 is concerned with a pre-limiting equation for the Stieltjes transform of the ESD (like (1.5), but before taking the limit $N\to\infty$). We improve the error term in the equation for using it further in the proof. In Sections 4 and 5 we present proofs of Theorems 2.1 and 2.2 respectively.
3 Equation for the Stieltjes transform
Proposition 3.1.
Let $W$ be defined in (1.2) and let the ESD $\nu$ of $S$ converge weakly in distribution to $\nu_{0}$ as $N\to\infty$. Then the Stieltjes transform $g_{\mu}(z)$ satisfies the following equation
$\operatorname{\mathbf{E}}g_{\mu}(z)-\operatorname{\mathbf{E}}g_{\nu}(z+\sigma^{2}\operatorname{\mathbf{E}}g_{\mu}(z))=O(N^{-1}),\quad N\to\infty.$ (3.1)
Proof.
Without loss of generality, we can assume that the matrix $S$ is diagonal. Indeed, let $S=U^{*}\Sigma U$, where $\Sigma$ is a diagonal matrix. Then the matrix
$UWU^{*}=\frac{1}{\sqrt{N}}URU^{*}+\Sigma$
has the same spectrum as $W$, and $URU^{*}$ has the same distribution as $R$. Thus, everywhere below in this proof we treat $S$ as a diagonal matrix.
$\operatorname{\mathbf{E}}\left\{g_{\mu}(z)\right\}=\dfrac{1}{N}\operatorname{\mathbf{E}}\left\{\operatorname{Tr}{G(z)}\right\}=\dfrac{1}{N}\sum\operatorname{\mathbf{E}}\left\{G_{ii}(z)\right\}$ (3.2)
where
$G(z)=(W-z)^{-1},\quad\Im{z}>0,$ (3.3)
is a resolvent of $W$.
Taking inverse of a block matrix, one obtains
$G_{ii}=\dfrac{1}{A_{i}},$ (3.4)
where
$A_{i}=W_{ii}-z-\bm{w}_{i}^{*}G^{(i)}\bm{w}_{i},$ (3.5)
$G^{(i)}=(W^{(i)}-z)^{-1},$ (3.6)
and $W^{(i)}$ is a $(N-1)\times(N-1)$ matrix obtained from $W$ by removing its $i$th column and its $i$th row, $\bm{w}_{i}$ is the $i$th column of $W$ without $W_{ii}$.
Since $S$ is diagonal, we have $\bm{w}_{i}=\frac{1}{\sqrt{N}}\bm{r}_{i}$. Let us transform the last term in (3.5)
$\begin{split}\bm{w}_{i}^{*}G^{(i)}\bm{w}_{i}=\frac{1}{N}\bm{r}_{i}^{*}G^{(i)}\bm{r}_{i}&=\dfrac{1}{N}\sum_{j\neq k;\ j,k\neq i}G_{jk}^{(i)}R_{ij}\overline{R}_{ik}+\dfrac{1}{N}\sum_{j=1}^{N}G_{jj}^{(i)}(|R_{jj}|^{2}-\sigma^{2})\\ &\quad{}+\sigma^{2}\left(\dfrac{1}{N}\operatorname{Tr}{G^{(i)}}-\operatorname{\mathbf{E}}\left\{\dfrac{1}{N}\operatorname{Tr}{G^{(i)}}\right\}\right)\\ &\quad{}+\sigma^{2}\operatorname{\mathbf{E}}\left\{\dfrac{1}{N}\operatorname{Tr}{G^{(i)}}-\dfrac{1}{N}\operatorname{Tr}{G}\right\}+\sigma^{2}\operatorname{\mathbf{E}}\left\{g_{\mu}(z)\right\}.\end{split}$ (3.7)
Denote
$r_{1}$ $=\dfrac{1}{N}\sum_{j\neq k;\ j,k\neq i}G_{jk}^{(i)}R_{ij}\overline{R}_{ik}$ (3.8)
$r_{2}$ $=\dfrac{1}{N}\sum_{j=1}^{N}G_{jj}^{(i)}(|R_{ij}|^{2}-\sigma^{2})$ (3.9)
$r_{3}$ $=\dfrac{\sigma^{2}}{N}\left(\operatorname{Tr}{G^{(i)}}-\operatorname{\mathbf{E}}\left\{\operatorname{Tr}{G^{(i)}}\right\}\right)$ (3.10)
$r_{4}$ $=\dfrac{\sigma^{2}}{N}\operatorname{\mathbf{E}}\left\{\operatorname{Tr}{G^{(i)}}-\operatorname{Tr}{G}\right\}.$ (3.11)
Thus we have
$A_{i}=A_{i}^{(0)}+\frac{1}{\sqrt{N}}R_{ii}+\sum\limits_{j=1}^{4}r_{j},$ (3.12)
where $A_{i}^{(0)}=S_{ii}-z-\sigma^{2}\operatorname{\mathbf{E}}\left\{g_{\mu}(z)\right\}$. We need the lemma
Lemma 3.2.
$\operatorname{\mathbf{E}}\left\{\left\lvert r_{j}\right\rvert^{2}\right\}\leq\frac{C}{N}$ for $j=1,2,3$, $\left\lvert r_{4}\right\rvert\leq\frac{C}{N}$.
The proof is given in Appendix.
$\operatorname{\mathbf{E}}\left\{g_{\mu}(z)-g_{\nu}(z+\sigma^{2}\operatorname{\mathbf{E}}\left\{g_{\mu}(z)\right\})\right\}=\operatorname{\mathbf{E}}\left\{\frac{1}{N}\sum\limits_{j=1}^{N}\frac{1}{A_{j}}-\frac{1}{N}\sum\limits_{j=1}^{N}\frac{1}{A_{j}^{(0)}}\right\}=\frac{1}{N}\sum\limits_{j=1}^{N}\operatorname{\mathbf{E}}\left\{\frac{1}{A_{j}}-\frac{1}{A_{j}^{(0)}}\right\}$ (3.13)
Estimate each term in (3.13)
$\operatorname{\mathbf{E}}\left\{\frac{1}{A_{j}}-\frac{1}{A_{j}^{(0)}}\right\}=\operatorname{\mathbf{E}}\left\{\frac{A_{j}^{(0)}-A_{j}}{\bigl{(}A_{j}^{(0)}\bigr{)}^{2}}\right\}+\operatorname{\mathbf{E}}\left\{\frac{(A_{j}^{(0)}-A_{j})^{2}}{\bigl{(}A_{j}^{(0)}\bigr{)}^{2}A_{j}}\right\}$ (3.14)
Consider the first term
$\left\lvert\operatorname{\mathbf{E}}\left\{\frac{A_{j}^{(0)}-A_{j}}{\bigl{(}A_{j}^{(0)}\bigr{)}^{2}}\right\}\right\rvert=\left\lvert\frac{r_{4}}{\bigl{(}A_{j}^{(0)}\bigr{)}^{2}}\right\rvert\leq\frac{\left\lvert r_{4}\right\rvert}{\left\lvert\Im z\right\rvert^{2}}\leq\frac{C}{N}.$ (3.15)
The second one
$\begin{split}\left\lvert\operatorname{\mathbf{E}}\left\{\frac{(A_{j}^{(0)}-A_{j})^{2}}{\bigl{(}A_{j}^{(0)}\bigr{)}^{2}A_{j}}\right\}\right\rvert&\leq\operatorname{\mathbf{E}}\left\{\left\lvert\frac{(A_{j}^{(0)}-A_{j})^{2}}{\bigl{(}A_{j}^{(0)}\bigr{)}^{2}A_{j}}\right\rvert\right\}\leq\frac{\operatorname{\mathbf{E}}\left\{\left\lvert A_{j}^{(0)}-A_{j}\right\rvert^{2}\right\}}{\left\lvert\Im z\right\rvert^{3}}\\ &\leq C\biggl{(}\frac{1}{N}\operatorname{\mathbf{E}}\left\{\left\lvert R_{jj}\right\rvert^{2}\right\}+\sum\limits_{i=1}^{4}\operatorname{\mathbf{E}}\left\{\left\lvert r_{i}\right\rvert^{2}\right\}\biggr{)}.\end{split}$ (3.16)
Using Lemma 3.2, one gets
$\begin{split}\left\lvert\operatorname{\mathbf{E}}\left\{\frac{(A_{j}^{(0)}-A_{j})^{2}}{\bigl{(}A_{j}^{(0)}\bigr{)}^{2}A_{j}}\right\}\right\rvert&\leq\frac{C}{N}.\end{split}$ (3.17)
Therefore, $\left\lvert\operatorname{\mathbf{E}}\left\{g_{\mu}(z)\right\}-\operatorname{\mathbf{E}}\left\{g_{\nu}(z+\sigma^{2}\operatorname{\mathbf{E}}\left\{g_{\mu}(z)\right\})\right\}\right\rvert\leq\frac{C}{N}$. ∎
Proof of Lemma 3.2.
Let $\operatorname{\mathbf{E}}_{i}$ be an expectation w.r.t. $\Re R_{ij}$ and $\Im R_{ij}$, $j=1,\dotsc,N$. We have
$\operatorname{\mathbf{E}}_{i}\left\{r_{1}\overline{r_{1}}\right\}=\dfrac{1}{N^{2}}\sum_{j_{1}\neq k_{1};\ j_{1},k_{1}\neq i}^{N}\sum_{j_{2}\neq k_{2};\ j_{2},k_{2}\neq i}^{N}G_{j_{1}k_{1}}^{(i)}\overline{G}_{j_{2}k_{2}}^{(i)}\operatorname{\mathbf{E}}_{i}\left\{R_{ij_{1}}\overline{R}_{ik_{1}}R_{ij_{2}}\overline{R}_{ik_{2}}\right\}.$ (3.18)
If $j_{1}\neq j_{2}$ and $j_{1}\neq k_{2}$, then $R_{ij_{1}}$ and $\overline{R}_{ik_{1}}R_{ij_{2}}\overline{R}_{ik_{2}}$ are independent, which implies
$\operatorname{\mathbf{E}}_{i}\left\{R_{ij_{1}}\overline{R}_{ik_{1}}R_{ij_{2}}\overline{R}_{ik_{2}}\right\}=\operatorname{\mathbf{E}}_{i}\left\{R_{ij_{1}}\right\}\operatorname{\mathbf{E}}_{i}\left\{\overline{R}_{ik_{1}}R_{ij_{2}}\overline{R}_{ik_{2}}\right\}=0.$ (3.19)
Similarly, the expectation is also zero if $k_{1}\neq j_{2}$ and $k_{1}\neq k_{2}$. Hence,
$\begin{split}\operatorname{\mathbf{E}}_{i}\left\{r_{1}\overline{r_{1}}\right\}&=\dfrac{1}{N^{2}}\sum_{\begin{subarray}{c}j_{1}\neq k_{1}\\ j_{1},k_{1}\neq i\end{subarray}}G_{j_{1}k_{1}}^{(i)}\overline{G}_{k_{1}j_{1}}^{(i)}\operatorname{\mathbf{E}}_{i}\left\{\left\lvert R_{ij_{1}}R_{ij_{2}}\right\rvert^{2}\right\}\\ &\quad{}+\dfrac{1}{N^{2}}\sum_{\begin{subarray}{c}j_{1}\neq k_{1}\\ j_{1},k_{1}\neq i\end{subarray}}^{N}\left\lvert G_{j_{1}k_{1}}^{(i)}\right\rvert^{2}\operatorname{\mathbf{E}}_{i}\left\{R_{ij_{1}}^{2}\overline{R}_{ik_{1}}^{2}\right\}\\ &\leq\frac{2}{N^{2}}\operatorname{Tr}G^{(i)}(G^{(i)})^{*}\leq\frac{2}{N}\left\lVert G^{(i)}\right\rVert^{2}\leq\frac{2}{N\left\lvert\Im z\right\rvert^{2}}.\end{split}$ (3.20)
So, $\operatorname{\mathbf{E}}\left\{\left\lvert r_{1}\right\rvert^{2}\right\}\leq\frac{C}{N}$. Let us proceed to derive an estimate on $r_{2}$. We get
$\begin{split}\operatorname{\mathbf{E}}_{i}\left\{r_{2}\overline{r_{2}}\right\}&=\dfrac{1}{N^{2}}\sum_{j_{1}\neq i}\sum_{j_{2}\neq i}G_{j_{1}j_{1}}^{(i)}\overline{G}_{j_{2}j_{2}}^{(i)}\operatorname{\mathbf{E}}_{i}\left\{(\left\lvert R_{ij_{1}}\right\rvert^{2}-\sigma^{2})(\left\lvert R_{ij_{2}}\right\rvert^{2}-\sigma^{2})\right\}\\ &=\dfrac{1}{N^{2}}\sum_{j\neq i}\left\lvert G_{jj}^{(i)}\right\rvert^{2}\operatorname{\mathbf{E}}_{i}\left\{(\left\lvert R_{ij}\right\rvert^{2}-\sigma^{2})^{2}\right\}\leq\frac{c_{4}}{N^{2}}\operatorname{Tr}G^{(i)}(G^{(i)})^{*}\leq\frac{c_{4}}{n\left\lvert\Im z\right\rvert^{2}},\end{split}$ (3.21)
because $\operatorname{\mathbf{E}}_{i}\left\{(\left\lvert R_{ij}\right\rvert^{2}-\sigma^{2})^{2}\right\}\leq c_{4}$. Further, $r_{3}$ can be written as follows
$\operatorname{\mathbf{E}}\left\{r_{3}\overline{r_{3}}\right\}=\operatorname{Var}\left\{\frac{1}{N}\operatorname{Tr}G^{(i)}\right\}\leq\frac{C}{N},$ (3.22)
where the last inequality is given in (18.2.13) of Theorem 18.2.3 in [6]. Since
$\left\lvert\operatorname{Tr}G^{(i)}-\operatorname{Tr}G\right\rvert\leq\frac{1}{\left\lvert\Im z\right\rvert}$, (3.23)
one has $r_{4}\leq\frac{C}{N}$. ∎
4 The limiting distribution of eigenvalues outside the bulk
In this section we derive the limiting ESD of outlayers. We will need the following lemma:
Lemma 4.1.
Functions $\Phi$ and $\omega_{\mu_{0}}$ defined in (1.9) and (1.6), respectively, satisfy the following: $\Phi(\omega_{\mu_{0}}(z))=z$ for any $z\in\mathbb{C}^{+}\setminus\operatorname{supp}{\mu_{0}}$.
Proof.
Use $\omega_{\mu_{0}}(z)$ (defined in (1.6)) and $\Phi(z)$ (defined in (1.9)) to compute $\Phi(\omega_{\mu_{0}}(z))$
$\Phi(\omega_{\mu_{0}}(z))=\omega_{\mu_{0}}(z)-\sigma^{2}g_{\nu_{0}}(\omega_{\mu_{0}}(z))$ (4.1)
Use right hand side of (1.6) for measure $\tau=\mu_{0}$ to get
$\Phi(\omega_{\mu_{0}}(z))=z+\sigma^{2}g_{\mu_{0}}(z)-\sigma^{2}g_{\nu_{0}}(\omega_{\mu_{0}}(z))$ (4.2)
Note that $g_{\mu_{0}}(z)=g_{\nu_{0}}(\omega_{\mu_{0}}(z))$ by (1.5). Thus, we show that for any $z\in\mathbb{C}^{+}\setminus\operatorname{supp}{\mu_{0}}$, $\Phi(\omega_{\mu_{0}}(z))=z$. ∎
Set
$\tilde{\nu}_{1}$ $=\frac{N}{r}(\nu-\nu_{0}),$ (4.3)
$\tilde{\mu}_{1}$ $=\frac{N}{r}(\mu-\mu_{0}).$ (4.4)
Lemma 4.2.
Assume that the signed measure $\tilde{\nu}_{1}$ weakly converges to a signed measure $\nu_{1}$. Then the signed measure $\tilde{\mu}_{1}$ also converges weakly in probability to a signed measure $\mu_{1}$ whose Stieltjes transform $g_{\mu_{1}}(z)$ satisfies
$g_{\mu_{1}}(z)=g_{\nu_{1}}(\omega_{\mu_{0}}(z))\omega^{\prime}_{\mu_{0}}(z)$ (4.5)
for any $z\in\mathbb{C}^{+}$.
Proof.
Note that (3.1) can be presented in the form
$f_{\mu}(z)=f_{\nu}(\operatorname{\mathbf{E}}\omega_{\mu}(z))+O(N^{-1}),$ (4.6)
where
$f_{\tau}(z):=\operatorname{\mathbf{E}}\left\{g_{\tau}(z)\right\}.$ (4.7)
Expand both sides of (4.6). Taking into account that
$\omega_{\mu}(z)=\omega_{\mu_{0}}(z)+\frac{r}{N}\sigma^{2}g_{\tilde{\mu}_{1}}(z),$
we get
$\begin{split}f_{\mu_{0}}(z)+\tfrac{r}{N}f_{\tilde{\mu}_{1}}(z)={}&f_{\nu_{0}}(\operatorname{\mathbf{E}}\omega_{\mu_{0}}(z))\\ &+\tfrac{r}{N}\Bigl{\{}f^{\prime}_{\nu_{0}}(\operatorname{\mathbf{E}}\omega_{\mu_{0}}(z))\sigma^{2}f_{\tilde{\mu}_{1}}(z)\\ &+f_{\tilde{\nu}_{1}}(\operatorname{\mathbf{E}}\omega_{\mu_{0}}(z))\Bigr{\}}+o\left(\tfrac{r}{N}\right).\end{split}$ (4.8)
Since $f_{\mu_{0}}(z)=f_{\nu_{0}}(\operatorname{\mathbf{E}}\omega_{\mu_{0}}(z))$, we have
$f_{\tilde{\mu}_{1}}(z)=f^{\prime}_{\nu_{0}}(\operatorname{\mathbf{E}}\omega_{\mu_{0}}(z))\sigma^{2}f_{\tilde{\mu}_{1}}(z)+f_{\tilde{\nu}_{1}}(\operatorname{\mathbf{E}}\omega_{\mu_{0}}(z))+o(1).$ (4.9)
Note that $\mu_{0}$ and $\nu_{0}$ are non-random measures, thus we can remove expected values from (4.9) and get
$f_{\tilde{\mu}_{1}}(z)=g^{\prime}_{\nu_{0}}(\omega_{\mu_{0}}(z))\sigma^{2}f_{\tilde{\mu}_{1}}(z)+f_{\tilde{\nu}_{1}}(\omega_{\mu_{0}}(z))+o(1).$ (4.10)
Solving (4.10) for $f_{\tilde{\mu}_{1}}(z)$, we obtain
$f_{\tilde{\mu}_{1}}(z)=f_{\tilde{\nu}_{1}}(\omega_{\mu_{0}}(z))\dfrac{1}{1-\sigma^{2}g^{\prime}_{\nu_{0}}(\omega_{\mu_{0}}(z))}+o(1).$ (4.11)
Now let us prove that the measure $\tilde{\mu}_{1}$ has a weak limit as $N\to\infty$. It is sufficient to check that
  1. (i)
    $\int_{M}^{\infty}d\tilde{\mu}_{1}$ uniformly in $N$, in probability goes to zero as $M\to\infty$;
  2. (ii)
    $\operatorname{Var}\{g_{\tilde{\mu}_{1}}(z)\}\to 0$ as $N\to\infty$;
  3. (iii)
    $f_{\tilde{\mu}_{1}}(z)$ converges uniformly on compacts as $N\to\infty$.
(iii) follows from Proposition 3.1. (ii) follows from the Poincaré inequality. Indeed, consider $g_{\tilde{\mu}_{1}}(z)$ as a function of $R$
$\Psi(R)=\Psi_{z}(R):=g_{\tilde{\mu}_{1}}(z)=\frac{1}{r}\operatorname{Tr}\left(N^{-1/2}R+S-z\right)^{-1}-\frac{N}{r}g_{\mu_{0}}(z).$ (4.12)
The Poincaré inequality implies
$\operatorname{Var}\{g_{\tilde{\mu}_{1}}(z)\}=\operatorname{Var}\{\Psi(R)\}\leq\operatorname{\mathbf{E}}\left\{\sigma^{2}\left\lvert\nabla\Psi(R)\right\rvert^{2}\right\}.$ (4.13)
Next, let us calculate the derivatives of $\Psi$. Differentiating the identity (Wz)G(z)=I(W-z)G(z)=I, where $G(z)$ is defined in (3.3), one gets
$\frac{\partial}{\partial R_{jk}}G(z)$ $=-N^{-1/2}G(z)(E_{jk}+E_{kj})G(z),\quad j<k,$ (4.14)
$\frac{\partial}{\partial R_{jj}}G(z)$ $=-N^{-1/2}G(z)E_{jj}G(z)$ (4.15)
with $E_{jk}$ be coordinate matrices, i.e. all their entries are zeros except one unit in the $j$th row and $k$th column. Taking trace of both sides and dividing by $r$, we have
$\frac{\partial}{\partial R_{jk}}\Psi(R)$ $=-\frac{1}{r\sqrt{N}}((G^{2})_{jk}+(G^{2})_{kj}),\quad j<k,$ (4.16)
$\frac{\partial}{\partial R_{jj}}\Psi(R)$ $=-\frac{1}{r\sqrt{N}}(G^{2})_{jj}.$ (4.17)
Therefore,
$\operatorname{\mathbf{E}}\left\{\sigma^{2}\left\lvert\nabla\Psi(R)\right\rvert^{2}\right\}\leq\frac{2\sigma^{2}}{r^{2}N}\operatorname{Tr}(G(z))^{4}\leq\frac{2\sigma^{2}}{r^{2}}\left\lVert(G(z))^{4}\right\rVert\leq\frac{2\sigma^{2}}{r^{2}\left\lvert\Im z\right\rvert^{4}}.$ (4.18)
Since $r\to\infty$ as $N\to\infty$, the combination of (4.13) and (4.18) yields (ii).
To prove (i), consider
$v_{2}:=\int t^{2}d\tilde{\mu}_{1}(t)=\frac{N}{r}\left(\int t^{2}d\mu(t)-\int t^{2}d\mu_{0}(t)\right)=\frac{N}{r}\left(\operatorname{Tr}W^{2}-\int t^{2}d\mu_{0}(t)\right).$ (4.19)
The second moment of the measure $\mu_{0}$ can be recovered from the equation on its Stieltjes transform (1.5)
$\int t^{2}d\mu_{0}(t)=\sigma^{2}+\int t^{2}d\nu_{0}(t).$ (4.20)
Then,
$\operatorname{Tr}W^{2}=\operatorname{Tr}\left(\frac{1}{\sqrt{N}}R+S\right)^{2}=\frac{1}{N}\operatorname{Tr}R^{2}+\frac{2}{\sqrt{N}}\operatorname{Tr}RS+\operatorname{Tr}S^{2}.$ (4.21)
Substituting (4.20) and (4.21) into (4.19), one obtains
$v_{2}=\frac{N}{r}\left[\frac{1}{N}\operatorname{Tr}R^{2}-\sigma^{2}\right]+\frac{2\sqrt{N}}{r}\operatorname{Tr}RS+\frac{N}{r}\left[\operatorname{Tr}S^{2}-\int t^{2}d\nu_{0}(t)\right]=:\Sigma_{1}+\Sigma_{2}+\Sigma_{3}.$ (4.22)
$\Sigma_{3}$ converges to $\int t^{2}d\nu_{1}$ by Assumption 2 and
$P(\left\lvert\Sigma_{i}\right\rvert>\varepsilon)\to 0,\quad i=1,2$ (4.23)
by the weak law of large numbers. Hence, as $N\to\infty$, $v_{2}$ converges to $\int t^{2}d\nu_{1}$. It implies (i).
Now we can take the limit in (4.11). We get
$f_{\mu_{1}}(z)=f_{\nu_{1}}(\omega_{\mu_{0}}(z))\dfrac{1}{1-\sigma^{2}g^{\prime}_{\nu_{0}}(\omega_{\mu_{0}}(z))}.$ (4.24)
Since $\Phi(\omega_{\mu_{0}}(z))=z$ for any $z\in\mathbb{C}^{+}\setminus\operatorname{supp}{\mu_{0}}$, where $\Phi$ is defined in (1.9), it is sufficient to verify that the denominator of (4.11) is equal to $\Phi^{\prime}(\omega_{\mu_{0}}(z))$, i.e.
$\Phi^{\prime}(\omega_{\mu_{0}}(z))=1-\sigma^{2}g^{\prime}_{\nu_{0}}(\omega_{\mu_{0}}(z)),$ (4.25)
Taking derivative in (1.9) we obtain
$\Phi^{\prime}(z)=1-\sigma^{2}g^{\prime}_{\nu_{0}}(z)$ (4.26)
Substituting $z\to\omega_{\mu_{0}}(z)$ into the equation above, we get (4.25).
By Lemma 4.1, $\Phi(\omega_{\mu_{0}}(z))=z$ for any $z\in\mathbb{C}^{+}\setminus\operatorname{supp}{\mu_{0}}$. Hence,
$\Phi^{\prime}(\omega_{\mu_{0}}(z))=\dfrac{1}{\omega^{\prime}_{\mu_{0}}(z)}.$ (4.27)
Use this fact and (4.11) to conclude that
$g_{\mu_{1}}(z)=g_{\nu_{1}}(\omega_{\mu_{0}}(z))\omega^{\prime}_{\mu_{0}}(z).$ (4.28)
Next, we present the proof of Theorem 2.1.
Proof of Theorem 2.1.
Let $z=x+iy$, $x\notin\operatorname{supp}\mu_{0}$.
As we have shown before in Lemma 4.2, the signed measure $\tilde{\mu}_{1}=\frac{N}{r}(\mu-\mu_{0})$ converges weakly in probability to the signed measure $\mu_{1}$ and its Stieltjes transform satisfies
$g_{\mu_{1}}(z)=g_{\nu_{1}}(\omega_{\mu_{0}}(z))\omega^{\prime}_{\mu_{0}}(z).$ (4.29)
Then,
$\begin{split}\rho_{\mu_{1}}(x)&=\lim_{y\to+0}\dfrac{1}{\pi}\Im g_{\mu_{1}}(x+iy)=\lim_{y\to+0}\dfrac{1}{\pi}\Im\left(g_{\nu_{1}}(\omega_{\mu_{0}}(x+iy))\omega^{\prime}_{\mu_{0}}(x+iy)\right)\\ &=\lim_{y\to+0}\dfrac{1}{\pi}\left(\Im g_{\nu_{1}}(\omega_{\mu_{0}}(x+iy))\Re\omega^{\prime}_{\mu_{0}}(x+iy)+\Re g_{\nu_{1}}(\omega_{\mu_{0}}(x+iy))\Im\omega^{\prime}_{\mu_{0}}(x+iy)\right)\end{split}$ (4.30)
The definition (1.6) implies that $\omega_{\mu_{0}}(z)$ is a holomorphic function for any $z\in\mathbb{C}\setminus\operatorname{supp}{\mu_{0}}$, thus it is continuous, and it’s derivative $\omega^{\prime}_{\mu_{0}}(z)$ is also continuous on the same set. Thus,
$\lim_{y\to+0}\Re\omega_{\mu_{0}}^{\prime}(x+iy)$ $=\Re\omega_{\mu_{0}}^{\prime}(x),$ (4.31)
$\lim_{y\to+0}\omega_{\mu_{0}}(x+iy)$ $=\omega_{\mu_{0}}(x)$ (4.32)
i.e., $\lim_{y\to+0}\Im\omega_{\mu_{0}}(x+iy)=0$ and $\lim_{y\to+0}\Re\omega_{\mu_{0}}(x+iy)=\omega_{\mu_{0}}(x)$. Changing variables $t=\Im\omega_{\mu_{0}}(x+iy)$, we get
$\lim_{y\to+0}\frac{1}{\pi}g_{\nu_{1}}(\omega_{\mu_{0}}(x+iy))=\lim_{t\to 0}\frac{1}{\pi}g_{\nu_{1}}(\omega_{\mu_{0}}(x)+o(t)+it).$ (4.33)
Let us check that $\omega_{\mu_{0}}(x+iy)$ has a positive imaginary part. From the definition (1.6) it follows that
$\Im\omega_{\mu_{0}}(x+iy)=\Im((x+iy)+\sigma^{2}g_{\mu_{0}}(x+iy))=y+\sigma^{2}\Im g_{\mu_{0}}(x+iy),$ (4.34)
By properties of the Stieltjes transform, $y\Im g_{\mu_{0}}(x+iy)>0$. Therefore, $\Im g_{\mu_{0}}(x+iy)>0$ for $y>0$. Also, for $x\notin\operatorname{supp}\mu_{0}$, there exists $\delta>0$ s.t. $(x-\delta,x+\delta)\cap\operatorname{supp}\mu_{0}=\emptyset$. We have for $0<y<\delta^{2}$
$\begin{split}\left\lvert\Im g_{\mu_{0}}(x+iy)\right\rvert=\left\lvert\int_{\left\lvert t-x\right\rvert\geq\delta}\frac{yd\mu_{0}(t)}{(t-x)^{2}+y^{2}}\right\rvert\leq C\int_{\left\lvert t-x\right\rvert\geq\delta}\frac{ydt}{(t-x)^{2}+y^{2}}\\ =C\int_{\left\lvert\tilde{t}\right\rvert\geq\delta/y}\frac{d\tilde{t}}{\tilde{t}^{2}+1}\leq C\frac{y}{\delta}<C\sqrt{y}.\ \end{split}$ (4.35)
Hence, $\Im\omega_{\mu_{0}}(x+iy)>0$, as a sum of two positive numbers, and
$\Im\omega_{\mu_{0}}(x+iy)\leq C\sqrt{y}\quad\text{as }y\to+0.$ (4.36)
Then, (4.33) transforms to
$\lim_{y\to+0}\frac{1}{\pi}\Im g_{\nu_{1}}(\omega_{\mu_{0}}(x+iy))=\lim_{t\to+0}\frac{1}{\pi}\Im g_{\nu_{1}}(\omega_{\mu_{0}}(x)+o(t)+it)=\rho_{\nu_{1}}(\omega_{\mu_{0}}(x)).$ (4.37)
Let us consider the second term in (4.30), $\Re g_{\nu_{1}}(\omega_{\mu_{0}}(x+iy))\Im\omega^{\prime}_{\mu_{0}}(x+iy)$. For $w=x_{1}+iy_{1}$ one gets
$\Re g_{\nu_{1}}(w)=\int\frac{(t-x_{1})d\nu_{1}(t)}{(t-x_{1})^{2}+y_{1}^{2}}$ (4.38)
and
$\begin{split}\left\lvert\Re g_{\nu_{1}}(w)\right\rvert\leq\left\lvert\int_{\left\lvert t-x_{1}\right\rvert\leq y_{1}}+\int_{\left\lvert t-x_{1}\right\rvert>y_{1}}\frac{(t-x_{1})d\nu_{1}(t)}{(t-x_{1})^{2}+y_{1}^{2}}\right\rvert\\ \leq\int_{\left\lvert t-x_{1}\right\rvert\leq y_{1}}\frac{d\left\lvert\nu_{1}\right\rvert(t)}{y_{1}}+\int_{\left\lvert t-x_{1}\right\rvert>y_{1}}\frac{d\left\lvert\nu_{1}\right\rvert(t)}{\left\lvert t-x_{1}\right\rvert}\leq C-C\log y_{1}.\end{split}$ (4.39)
Thus, taking into account (4.36), one obtains, as $y\to+0$,
$\left\lvert\Re g_{\nu_{1}}(\omega_{\mu_{0}}(x+iy))\right\rvert\leq C(1-\log\Im\omega_{\mu_{0}}(x+iy))\leq C(1-\log y).$ (4.40)
Next,
$\Im\omega^{\prime}_{\mu_{0}}(x+iy)=\sigma^{2}\Im g_{\mu_{0}}^{\prime}(x+iy)=2\sigma^{2}\int\frac{(t-x)yd\mu_{0}(t)}{((t-x)^{2}+y^{2})^{2}}.$ (4.41)
It implies
$\lim\limits_{y\to+0}\frac{\Im\omega^{\prime}_{\mu_{0}}(x+iy)}{y}=\sigma^{2}\int\frac{d\mu_{0}(t)}{(t-x)^{3}}.$ (4.42)
Hence, $\Im\omega^{\prime}_{\mu_{0}}(x+iy)\leq Cy$ and
$\left\lvert\Re g_{\nu_{1}}(\omega_{\mu_{0}}(x+iy))\Im\omega^{\prime}_{\mu_{0}}(x+iy)\right\rvert\leq C(1-\log y)y\to 0\quad\text{as }y\to+0.$ (4.43)
A combination of (4.30), (4.37), and (4.43) gives us
$\rho_{\mu_{1}}(x)=\rho_{\nu_{1}}(\omega_{\mu_{0}}(x))\omega_{\mu_{0}}^{\prime}(x).$ (4.44)
Finally,
$\mu_{1}(\Delta)=\int_{\Delta}\rho_{\mu_{1}}(x)dx=\int_{\Delta}\rho_{\nu_{1}}(\omega_{\mu_{0}}(x))\omega_{\mu_{0}}^{\prime}(x)dx=\int_{\omega_{\mu_{0}}(\Delta)}\rho_{\nu_{1}}(x_{1})dx_{1}=\nu_{1}(\omega_{\mu_{0}}(\Delta)).$ (4.45)
5 Proof of Theorem 2.2
In this section, we present the proof of Theorem 2.2.
According to Theorem 2.1, $\operatorname{supp}\mu_{1}^{+}=[\Phi(a),\Phi(b)]$. Pick an arbitrary positive integer $k$. Let the points $\zeta_{1}\leq\dotsb\leq\zeta_{k-1}$ split the segment $\Phi(a),\Phi(b)]$ into $k$ equal parts, i.e.
$\zeta_{i}=\Phi(a)+\frac{i}{k}(\Phi(b)-\Phi(a)),\quad 0<i<k.$ (5.1)
Continuity of $\Phi(x)$ and (2.4) imply $\lim_{N\to\infty}\max_{1\leq j\leq r}\operatorname{dist}(\lambda_{j}(W),[\Phi(a),\Phi(b)])=0$ and $\lambda_{1}(W)\leq c$ for some $c$. Fix arbitrary $\varepsilon_{1},\varepsilon_{2}>0$ and consider functions
$h_{i}(x)=\begin{cases}0,&x<\zeta_{i}\text{ or }x>c+\varepsilon_{1}\\ \frac{x-\zeta_{i}}{\varepsilon_{1}},&\zeta_{i}\leq x<\zeta_{i}+\varepsilon_{1}\\ 1,&\zeta_{i}+\varepsilon_{1}\leq x<c\\ \frac{c+\varepsilon_{1}-x}{\varepsilon_{1}},&c\leq x\leq c+\varepsilon_{1}\end{cases}$ (5.2)
[…]
Take any positive $\varepsilon_{3}$ such that
$\varepsilon_{3}<\frac{1}{2}\min_{i}\int(h_{i}(x)-h_{i+1}(x))d\mu_{1}(x).$ (5.3)
Assumption 3 and Theorem 2.1 yield that there exists $N_{0}$ such that for any $N>N_{0}$
$\sum_{i}\left\lvert\tfrac{N}{r}\int h_{i}(\Phi(x))d(\nu(x)-\nu_{0}(x))-\int h_{i}(\Phi(x))d\nu_{1}(x)\right\rvert<\varepsilon_{3},$ (5.4)
$P\left(\sum_{i}\left\lvert\tfrac{N}{r}\int h_{i}(x)d\mu(x)-\int h_{i}(x)d\mu_{1}(x)\right\rvert>\varepsilon_{3}\right)<\varepsilon_{2}.$ (5.5)
Fix some $N>N_{0}$ and let $\lambda_{j(N)}(W)\in[\zeta_{m},\zeta_{m+1}]$. Using (5.4), we get
$\begin{split}\#\{{i\mid\lambda_{i}(S)\geq\omega_{\mu_{0}}(\zeta_{m+3})}\}&=N\int 1_{\omega_{\mu_{0}}(\zeta_{m+3})\leq x\leq c}(x)d\nu(x)\leq N\int h_{m+2}(\Phi(x))d\nu(x)\\ &<r\int h_{m+2}(\Phi(x))d\nu_{1}(x)+r\varepsilon_{3}.\end{split}$ (5.6)
By Theorem 2.1 and by the choice of $\varepsilon_{3}$ (5.3),
$\int h_{m+2}(\Phi(x))d\nu_{1}(x)=\int h_{m+2}(x)d\mu_{1}(x)<\int h_{m+1}(x)d\mu_{1}(x)-2\varepsilon_{3}$ (5.7)
(5.5) implies
$\begin{split}r\int h_{m+1}(x)d\mu_{1}(x)-r\varepsilon_{3}<N\int h_{m+1}(x)d\mu(x)\leq\#\{{i\mid\lambda_{i}(W)>\zeta_{m+1}}\}<j(N)\end{split}$ (5.8)
with probability greater or equal to $1-\varepsilon_{2}$. Hence, combining (5.6)–(5.8) one obtains $\lambda_{j(N)}(S)<\omega_{\mu_{0}}(\zeta_{m+3})$. Similarly,
$\begin{split}\#\{{i\mid\lambda_{i}(S)\geq\omega_{\mu_{0}}(\zeta_{m-2})}\}&\geq N\int h_{m-2}(\Phi(x))d\nu(x)>r\int h_{m-2}(\Phi(x))d\nu_{1}(x)-r\varepsilon_{3}\\ &>r\int h_{m-1}(x)d\mu_{1}(x)+r\varepsilon_{3}>N\int h_{m-1}(x)d\mu(x)\\ &\geq\#\{{i\mid\lambda_{i}(W)>\zeta_{m}}\}\geq j(N)\end{split}$ (5.9)
and $\omega_{\mu_{0}}(\zeta_{m-2})<\lambda_{j(N)}(S)$. Thus, $\zeta_{m-2}<\Phi(\lambda_{j(N)}(S))<\zeta_{m+3}$ which implies
$P\left[\left\lvert\lambda_{j(N)}(W)-\Phi(\lambda_{j(N)}(S))\right\rvert<\frac{3}{k}(\Phi(b)-\Phi(a))\right]>1-\varepsilon_{2}.$ (5.10)
Since $k$ and $\varepsilon_{2}$ are arbitrary, the last equation yields the convergence (2.5) in probability.
Appendix A Numerical simulations
For the numerical simulations below, the following DNN was considered:
Networks were trained for 25 epochs on the Fashion MNIST dataset. Each network has 3 weight layers of the following sizes: $784\times N$, $N\times N$, $N\times 10$, where the matrix size $N$ changes from 200 to 3000, with step 200. For every $N$ the network achieves test accuracy 88-89%.
Figure 2: Numerical simulations of a DNN with 3 layers. Two figures correspond to two different realizations with the same architecture.
These numerical simulations demonstrate the dependence of the number of eigenvalues of the ‘‘signal’’ matrix on the size of the middle layer matrix.
Acknowledgement. The work of I.A. was supported by the Grant “International Multilateral Partnerships for Resilient Education and Science System in Ukraine” IMPRESS-U: N7114 funded by US National Academy of Science and Office of Naval Research. The work of L.B. and M.K. was partially supported by the NSF Grant IMPRESS-U: N2401227. The authors are grateful to Prof. Mariya Shcherbina and Prof. Tatyana Shcherbina for fruitful discussions and useful suggestions.