### v3

# Asymptotic behavior of eigenvalues of large rank perturbations of large random matrices

Abstract
The paper is concerned with deformed Wigner random matrices. These matrices are closely related to Deep Neural Networks (DNNs): weight matrices of trained DNNs could be represented in the form $R+S$, where $R$ is random and $S$ is highly correlated. The spectrum of such matrices plays a key role in rigorous underpinning of the novel pruning technique based on Random Matrix Theory. In practice, the spectrum of the matrix $S$ can be rather complicated. In this paper, we develop an asymptotic analysis for the case of full rank $S$ with increasing number of outlier eigenvalues.
1 Introduction
Random Matrix Theory (RMT) is a classical theory that has been developing for more than 70 years. Initially, RMT arose from problems in nuclear physics and found applications in mathematics, physics, finance, and many other disciplines. Recently, new problems have been arising in the area of Machine Learning. Indeed, often the weight matrices of Deep Neural Networks (DNNs) are initialized randomly. Moreover, modern DNNs have large weight matrices, which is why their spectral properties can be described by the asymptotic behavior of $N\times N$ random matrices as $N$ goes to infinity.
1.1 Formulation of the problem
Martin and Mahoney in [6] showed that weight matrices of trained DNNs can be modeled by large $N\times M$ random matrices of the form
$W=\frac{1}{\sqrt{N}}R+S,$ (1.1)
where $R$ is a rectangular matrix whose entries are i.i.d. random variables with zero mean and variance $\sigma^{2}$ and $S$ is a real non-random matrix (or random, but strongly correlated).
The present paper is concerned with a similar but simpler $N\times N$ deformed Wigner Ensemble matrices of the form
$W=\frac{1}{\sqrt{N}}R+S,$ (1.2)
where $R$ is restricted to be a square real symmetric matrix whose entries are i.i.d. random variables with zero mean and variance
$\operatorname{\mathbf{E}}\left\{R_{jk}^{2}\right\}=(1+\delta_{jk})\sigma^{2},$ (1.3)
$\operatorname{\mathbf{E}}$ denotes the expectation and $\delta_{jk}$ is $1$ if $j=k$ and 0 otherwise. $S$ is a real symmetric non-random matrix.
The empirical spectral distribution (ESD, a.k.a. the Normalized Counting Measure) for an $N\times N$ symmetric matrix $X$ is defined as:
$\mu_{N}(\Delta)=\frac{1}{N}\sum_{j=1}^{N}\mathbbm{1}_{\Delta}(\lambda_{j}(X))$ (1.4)
where $\Delta$ is an interval in $\mathbb{R}$,
$\lambda_{1}(X)\geq\lambda_{2}(X)\geq\dotsb\geq\lambda_{N}(X)$ (1.5)
are eigenvalues of $X$ and $\mathbbm{1}_{\Delta}$ is a characteristic function of the set $\Delta$. For a probability measure $\tau$ on $\mathbb{R}$, denote by $g_{\tau}$ its Stieltjes transform defined for $z\in\mathbb{C}\setminus\operatorname{supp}(\tau)$ by
$g_{\tau}(z)=\int\limits_{\mathbb{R}}\frac{d\tau(t)}{t-z}.$ (1.6)
In 1972 Pastur established that the ESD $\mu$ of the matrices (1.2) converges as $N\to\infty$, assuming that the ESD $\nu$ of $S$ converges to the measure $\nu_{0}$ [7], see also [1]. It was also shown that the Stieltjes transform $g_{\mu_{0}}(z)$ of the limiting ESD $\mu_{0}$ of the matrix $W$ defined in (1.2) satisfies the equation
$g_{\mu_{0}}(z)=g_{\nu_{0}}(\omega_{\mu_{0}}(z)),\quad z\in\mathbb{C}_{+},$ (1.7)
where for any measure $\tau$ we denote
$\omega_{\tau}(z)=z+\sigma^{2}g_{\tau}(z).$ (1.8)
In the particular case of $\nu_{0}$ being a delta-measure at zero, $\mu_{0}$ coincides with a well-known semicircular law $\mu_{sc}$ [10]
$\frac{d\mu_{sc}}{dx}(x)=\frac{1}{2\pi\sigma^{2}}\sqrt{4\sigma^{2}-x^{2}}\mathbbm{1}_{[-2\sigma,2\sigma]}(x).$ (1.9)
The individual asymptotic behavior of the eigenvalues $\lambda_{j}(W)$ depends on the individual asymptotic behavior of the eigenvalues $\lambda_{j}(S)$. The eigenvalues of $S$ may either converge to the bulk (i.e., $\operatorname{supp}\nu_{0}$) or not. Let us assume that the largest $r(N)$ eigenvalues $\lambda_{j}(S)$ do not converge to the bulk, but that the other do.
$\max_{1\leq j\leq r(N)}\operatorname{dist}(\lambda_{j}(S),\operatorname{supp}\nu_{0})>\varepsilon>0;$ (1.10)
$\lim_{N\to\infty}\max_{r(N)<j\leq N}\operatorname{dist}(\lambda_{j}(S),\operatorname{supp}\nu_{0})=0.$ (1.11)
The eigenvalues satisfying the property (1.10) are called spikes or outliers.
For $R$ being a matrix from a Gaussian Unitary Ensemble and for a low rank matrix $S$ ($\lambda_{j}(S)=0$ for $j>r(N)$) with $\lim\limits_{N\to\infty}\frac{r(N)}{N}=0$, Péché established the behavior of the largest eigenvalue $\lambda_{1}(W)$ in [9]. In particular, for fixed $r(N)=r^{\prime}$ it was shown that if $\lim\limits_{N\to\infty}\lambda_{1}(S)=\theta_{1}$ then almost surely
$\lim\limits_{N\to\infty}\lambda_{1}(W)=\begin{cases}2\sigma,\quad\text{if }\theta_{1}<\sigma,\\ \theta_{1}+\frac{\sigma^{2}}{\theta_{1}},\quad\text{if }\theta_{1}\geq\sigma.\end{cases}$ (1.12)
The general case where $R$ is drawn from the Wigner Ensemble was considered in [3] and [4]. In order to describe these results, let us introduce a function
$\Phi(z)=z-\sigma^{2}g_{\nu_{0}}(z).$ (1.13)
Note that if $\nu_{0}$ is a delta-function at zero, then $\Phi(\theta_{1})$ coincides with the second line in (1.12). Capitaine et al [3] dealt with fixed $r(N)=r^{\prime}$ and with $S$ of general form. They showed the following. Let $\lim\limits_{N\to\infty}\lambda_{j}(S)=\theta_{j}$, $j=1,\dotsc,r^{\prime}$ and let the common distribution of the entries of $R$ be symmetric and satisfy a Poincaré inequality. Then $\lambda_{j}(W)$ converges almost surely to $\Phi(\theta_{j})$ if $\Phi^{\prime}(\theta_{j})>0$, and to the bulk otherwise. In [4] Huang showed a similar convergence of eigenvalues, but in the case where $r(N)\to\infty$, $r(N)\ll N$, and $S$ is a diagonal low rank matrix, namely, all the eigenvalues of $S$ except the spikes are zeros.
Let $\mu$ and $\nu$ be the ESDs of $W$ and $S$, respectively, and let $S$ have exactly $r=r(N)$ non-zero eigenvalues. Denote
$\eta_{N}(\Delta)=\frac{1}{r}\sum_{\lambda_{j}(S)\neq 0}\mathbbm{1}_{\Delta}(\lambda_{j}(S)).$ (1.14)
We impose the following assumptions on the matrix $S$ and its spectrum.
Assumption 1.
As $N\to\infty$, the ESD $\nu$ weakly converges to the measure $\nu_{0}$.
Assumption 2.
There are $r=r(N)$ eigenvalues of $S$ outside $\operatorname{supp}\nu_{0}$. Let $\lim\limits_{N\to\infty}r(N)=\infty$ and $r(N)=o(N)$ as $N\to\infty$.
Assumption 3.
As $N\to\infty$, the measure $\frac{N}{r}(\nu-\nu_{0})$ weakly converges to the signed measure $\nu_{1}$, $\nu_{1}(\mathbb{R}\setminus\operatorname{supp}\nu_{0})=-\nu_{1}(\operatorname{supp}\nu_{0})=1$. Moreover, $\lim\limits_{N\to\infty}\frac{N}{r}\int t^{2}d(\nu-\nu_{0})(t)=\int t^{2}d\nu_{1}(t)$.
The crucial distinction from the previous results is that we simultaneously consider the general form of $\nu_{0}$ and $r(N)\to\infty$. In [3] the measure $\nu_{0}$ is of general form, but the deterministic matrix $S$ has a finite rank $r(N)=r^{\prime}$ independent of $N$, while in [4] $r(N)$ goes to infinity but $\nu_{0}$ is a delta-measure at zero.
1.2 Motivation: weight matrices of DNNs
DNNs is a broad class of computational algorithms inspired by the human brain. They are used for classification, approximation, and image recognition, among others. DNNs have become central to modern machine learning, and understanding their mathematical properties remains an important challenge.
Let us consider a DNN to solve a classification problem. The setup of a classification problem is as follows. There is a set of some objects $\mathcal{S}$. Each object $s\in\mathcal{S}$ belongs to one of $K$ classes. Given the object, one wants to determine its class. We search for a solution in the form of a function $\varphi\colon\mathcal{S}\to[0;1]^{K}$, interpreting the components of $\varphi(s)$ as probabilities that $s$ belongs to each class.
From the mathematical point of view, a feedforward fully-connected DNN for solving the classification problem is a function of the following form:
$\varphi(\cdot,\alpha)=\sigma\circ\lambda\circ f_{M}(\cdot,\alpha^{M})\circ\lambda\circ f_{M-1}(\cdot,\alpha^{M-1})\circ\cdots\circ\lambda\circ f_{1}(\cdot,\alpha^{1})$ (1.15)
where each $f_{k}:\mathbb{R}^{N_{k-1}}\to\mathbb{R}^{N_{k}}$, $f_{k}(x,\alpha^{k})=W_{k}x+\beta_{k}$ is an affine function, $W_{k}$ is an $N_{k}\times N_{k-1}$ weight matrix, $\beta_{k}\in\mathbb{R}^{N_{k}}$ is a bias vector, $\lambda:\mathbb{R}\to\mathbb{R}$ is a nonlinear activation function, which is applied component-wise (for example, absolute value), $\sigma$ is a function which normalizes vector’s components to probabilities. The entries of $W_{k}$ and the biases $\beta_{k}$ are the parameters $\alpha^{k}$ of a DNN. The parameters are optimized by minimizing the loss function, using, for example, stochastic gradient descent. See [2] for more details.
Pruning parameters of a DNN is the process of removing certain weights (set them equal to zero) from the network in a way which preserves the overall performance. The goal is to eliminate components that contribute little to the output of the DNN function, resulting in a more sparse and more efficient architecture. Pruning parameters of deep neural networks is an important task because it reduces the complexity of the function, leading to smaller, faster, and more efficient models. This not only decreases memory and computational requirements — crucial for deployment on resource-constrained devices — but can also improve generalization by preventing overfitting. In [5] Berlyand et al. provided mathematical underpinning to the pruning algorithm in the following setting: the weight matrix to be pruned can be represented as a sum of random and low-rank (‘‘signal’’) components with non-intersecting spectra (1.1). They call it Marchenko-Pastur(MP)- pruning as the key step is setting the singular values of the weight matrix smaller than the right edge of the support of Marchenko-Pastur distribution to zero, which helps remove unnecessary randomness from the DNN. The asymptotical behavior of the spectrum of matrices (1.1) plays a key role.
The fixed low-rank structure of the ‘‘signal’’ matrix allows certain analytical tools to be employed, leading to provable guaranties on network behavior, particularly the bounds on accuracy and loss after pruning. However, numerical experiments reveal that this low-rank assumption does not always hold in practice. Realistic weight matrices often have a spectrum that has a ‘‘signal’’ part with the number of non-zero eigenvalues increasing with respect to the matrix size (see Figure 1). See Appendix A for more details. These observations suggest that the existing theoretical results, while insightful, are not fully representative of the behavior seen in real-world networks.
Figure 1: Numeric simulation of a DNN with 3 layers. Show the dependence of the number of outlier eigenvalues of the ‘‘signal’’ matrix, so called spikes, on the size of the matrix.
In this work, we study a mathematical problem that is applicable in a simplified yet insightful setting: a DNN with two consecutive layers of equal size. This structure implies that the weight matrix connecting these layers is square. Furthermore, we impose an additional assumption of symmetry on the weight matrix, which is not applicable for DNNs directly, but simplifies the theoretical analysis.
Motivated by the gap between existing rigorous proofs and numerical results, our goal is to generalize the existing theoretical framework to accommodate symmetric weight matrices without a low-rank constraint or a fixed number of spikes. By relaxing these assumptions, we aim to provide a more comprehensive understanding of the accuracy of the network and the behavior of the loss function. This generalization brings the theory closer to practical scenarios, namely the ‘‘bulk decay’’ case (see [6, Figure 12(b) and 12(d)]), where the distribution of eigenvalues looks not like the MP-distribution plus some spikes, but like the dilated MP distribution plus some spikes. Moreover, this generalization raises interesting mathematical questions about the behavior of random matrices beyond the low-rank perturbation setting. Through this work, we seek to bridge the divide between theory and numerical computations, contributing to a more robust foundation for analyzing deep learning algorithms.
2 Main results
In this section, we present our main results. We first consider the measure $\frac{N}{r}(\mu-\mu_{0})$, which can be interpreted as an ESD of eigenvalues outside the bulk. The following theorem proves the existence of the limiting measure and describes it.
Theorem 2.1.
Let $W$ be a random matrix of the form (1.2). Let the third and fourth moments of the entries of $R$ be finite. Let $\mu$ and $\nu$ be the ESDs of the matrices $W$ and $S$, respectively, $\mu_{0}$ be the limit of $\mu$ as $N\to\infty$, and let Assumptions 1–3 hold. Then a signed measure $\frac{N}{r}(\mu-\mu_{0})$ converges weakly in distribution, as $N\to\infty$, to a non-random measure $\mu_{1}$. Moreover, if $\mu_{1}$ has a density, then for any measurable set $\Delta$ that does not intersect with $\operatorname{supp}\mu_{0}$
$\mu_{1}(\Delta)=\nu_{1}(\omega_{\mu_{0}}(\Delta)).$ (2.1)
The next theorem describes the asymptotic behavior of individual eigenvalues. To introduce the theorem, we need more notation. For any matrix $X$, let its eigenvalues be
$\lambda_{1}(X)\geq\lambda_{2}(X)\geq\dotsb\geq\lambda_{N}(X).$ (2.2)
Also, denote the positive part of $\nu_{1}$ by $\nu_{1}^{+}$.
Theorem 2.2.
Let $W$ be a random matrix of the form (1.2). Let the third and fourth moments of the entries of $R$ be finite. Let $\mu$ and $\nu$ be the ESDs of matrices $W$ and $S$, respectively, and let Assumptions 1–3 hold. Additionally, we assume that $\operatorname{supp}\nu_{1}^{+}=[\omega_{\mu_{0}}(a),\omega_{\mu_{0}}(b)]$, $\operatorname{dist}(\omega_{\mu_{0}}(a),\operatorname{supp}\nu_{0})>0$, $\lambda_{j}(S)$, $1\leq j\leq r$, are outside $\operatorname{supp}\nu_{0}$ and
$\lim_{N\to\infty}\max_{1\leq j\leq N}\operatorname{dist}(\lambda_{j}(S),\operatorname{supp}\nu_{0}\cup[\omega_{\mu_{0}}(a),\omega_{\mu_{0}}(b)])=0;$ (2.3)
$\lim_{N\to\infty}\max_{1\leq j\leq r}\operatorname{dist}(\lambda_{j}(S),[\omega_{\mu_{0}}(a),\omega_{\mu_{0}}(b)])=0.$ (2.4)
Then for any $j(N)\leq r(N)$
$\lambda_{j(N)}(W)-\Phi(\lambda_{j(N)}(S))\to 0,\text{ as }N\to\infty$ (2.5)
in probability, where $\Phi(x)$ is defined in (1.13).
Remarks
  1. 1.
    If the sequence $j(N)$ is chosen such that $\lim_{N\to\infty}\lambda_{j(N)}(S)=\theta$, then $\lambda_{j(N)}(W)\to\Phi(\theta)$, as $N\to\infty$.
  2. 2.
    Theorem 2.2 agrees well with the results obtained in [3, 4, 9].
The paper is organized as follows. Section 3 is concerned with a pre-limiting equation for the Stieltjes transform of the ESD (like (1.7), but before taking the limit $N\to\infty$). We improve the error term in the equation to use it further in the proof. In Sections 4 and 5 we present proofs of Theorems 2.1 and 2.2, respectively.
3 Equation for the Stieltjes transform
In this section, we prove that an error term in the pre-limiting equation for the Stieltjes transform of the ESD has the order $O(N^{-1})$. We do it in two steps: first, for the case of deformed Gaussian Orthogonal Ensemble (GOE)
$\tilde{W}=N^{-1/2}H+S,$ (3.1)
where $N^{-1/2}H$ is drawn from GOE, and then in the general case. Proposition 3.1 covers the first step.
Proposition 3.1.
Let $\tilde{W}$ be defined in (3.1). Let the ESD $\nu$ of $S$ converge weakly to $\nu_{0}$ as $N\to\infty$. Then the Stieltjes transform $g_{\mu}(z)$ satisfies the following equation
$\operatorname{\mathbf{E}}g_{\mu}(z)-\operatorname{\mathbf{E}}g_{\nu}(z+\sigma^{2}\operatorname{\mathbf{E}}g_{\mu}(z))=O(N^{-1}),\quad N\to\infty.$ (3.2)
Proof.
Without loss of generality, we can assume that the matrix $S$ is diagonal. Indeed, Let $S=U^{*}\Sigma U$, where $\Sigma$ is a diagonal matrix. Then the matrix
$U\tilde{W}U^{*}=\frac{1}{\sqrt{N}}UHU^{*}+\Sigma$
has the same spectrum as $\tilde{W}$, and $UHU^{*}$ has the same distribution as $H$. Thus, it is enough to prove the claim for the matrices $W$ of the form
$W=N^{-1/2}H+\Sigma.$ (3.3)
In a standard way, we write the Stieltjes transform in terms of the resolvent of $W$
$\operatorname{\mathbf{E}}\left\{g_{\mu}(z)\right\}=\dfrac{1}{N}\operatorname{\mathbf{E}}\left\{\operatorname{Tr}{G(z)}\right\}=\dfrac{1}{N}\sum\operatorname{\mathbf{E}}\left\{G_{ii}(z)\right\}$ (3.4)
where
$G(z)=(W-z)^{-1},\quad\Im{z}>0,$ (3.5)
is a resolvent of $W$.
Taking the inverse of a block matrix, one obtains
$G_{ii}=\dfrac{1}{A_{i}},$ (3.6)
where
$A_{i}=W_{ii}-z-\bm{w}_{i}^{*}G^{(i)}\bm{w}_{i},$ (3.7)
$G^{(i)}=(W^{(i)}-z)^{-1},$ (3.8)
and $W^{(i)}$ is an $(N-1)\times(N-1)$ matrix obtained from $W$ by removing its $i$th column and its $i$th row, $\bm{w}_{i}$ is the $i$th column of $W$ without $W_{ii}$.
Since $\Sigma$ is diagonal, we have $\bm{w}_{i}=\frac{1}{\sqrt{N}}\bm{r}_{i}$. Let us transform the last term in (3.7)
$\begin{split}\bm{w}_{i}^{*}G^{(i)}\bm{w}_{i}=\frac{1}{N}\bm{r}_{i}^{*}G^{(i)}\bm{r}_{i}&=\dfrac{1}{N}\sum_{j\neq k;\ j,k\neq i}G_{jk}^{(i)}R_{ij}\overline{R}_{ik}+\dfrac{1}{N}\sum_{j=1}^{N}G_{jj}^{(i)}(|R_{jj}|^{2}-\sigma^{2})\\ &\quad{}+\sigma^{2}\left(\dfrac{1}{N}\operatorname{Tr}{G^{(i)}}-\operatorname{\mathbf{E}}\left\{\dfrac{1}{N}\operatorname{Tr}{G^{(i)}}\right\}\right)\\ &\quad{}+\sigma^{2}\operatorname{\mathbf{E}}\left\{\dfrac{1}{N}\operatorname{Tr}{G^{(i)}}-\dfrac{1}{N}\operatorname{Tr}{G}\right\}+\sigma^{2}\operatorname{\mathbf{E}}\left\{g_{\mu}(z)\right\}.\end{split}$ (3.9)
Denote
$r_{1}$ $=\dfrac{1}{N}\sum_{j\neq k;\ j,k\neq i}G_{jk}^{(i)}R_{ij}\overline{R}_{ik}$ (3.10)
$r_{2}$ $=\dfrac{1}{N}\sum_{j=1}^{N}G_{jj}^{(i)}(|R_{ij}|^{2}-\sigma^{2})$ (3.11)
$r_{3}$ $=\dfrac{\sigma^{2}}{N}\left(\operatorname{Tr}{G^{(i)}}-\operatorname{\mathbf{E}}\left\{\operatorname{Tr}{G^{(i)}}\right\}\right)$ (3.12)
$r_{4}$ $=\dfrac{\sigma^{2}}{N}\operatorname{\mathbf{E}}\left\{\operatorname{Tr}{G^{(i)}}-\operatorname{Tr}{G}\right\}.$ (3.13)
Thus, we have
$A_{i}=A_{i}^{(0)}+\frac{1}{\sqrt{N}}R_{ii}+\sum\limits_{j=1}^{4}r_{j},$ (3.14)
where $A_{i}^{(0)}=S_{ii}-z-\sigma^{2}\operatorname{\mathbf{E}}\left\{g_{\mu}(z)\right\}$. We need the lemma
Lemma 3.2.
$\operatorname{\mathbf{E}}\left\{\left\lvert r_{j}\right\rvert^{2}\right\}\leq\frac{C}{N}$ for $j=1,2,3$, $\left\lvert r_{4}\right\rvert\leq\frac{C}{N}$.
The proof of the lemma is given in Appendix B.
$\operatorname{\mathbf{E}}\left\{g_{\mu}(z)-g_{\nu}(z+\sigma^{2}\operatorname{\mathbf{E}}\left\{g_{\mu}(z)\right\})\right\}=\operatorname{\mathbf{E}}\left\{\frac{1}{N}\sum\limits_{j=1}^{N}\frac{1}{A_{j}}-\frac{1}{N}\sum\limits_{j=1}^{N}\frac{1}{A_{j}^{(0)}}\right\}=\frac{1}{N}\sum\limits_{j=1}^{N}\operatorname{\mathbf{E}}\left\{\frac{1}{A_{j}}-\frac{1}{A_{j}^{(0)}}\right\}$ (3.15)
Estimate each term in (3.15)
$\operatorname{\mathbf{E}}\left\{\frac{1}{A_{j}}-\frac{1}{A_{j}^{(0)}}\right\}=\operatorname{\mathbf{E}}\left\{\frac{A_{j}^{(0)}-A_{j}}{\bigl(A_{j}^{(0)}\bigr)^{2}}\right\}+\operatorname{\mathbf{E}}\left\{\frac{(A_{j}^{(0)}-A_{j})^{2}}{\bigl(A_{j}^{(0)}\bigr)^{2}A_{j}}\right\}$ (3.16)
Consider the first term
$\left\lvert\operatorname{\mathbf{E}}\left\{\frac{A_{j}^{(0)}-A_{j}}{\bigl(A_{j}^{(0)}\bigr)^{2}}\right\}\right\rvert=\left\lvert\frac{r_{4}}{\bigl(A_{j}^{(0)}\bigr)^{2}}\right\rvert\leq\frac{\left\lvert r_{4}\right\rvert}{\left\lvert\Im z\right\rvert^{2}}\leq\frac{C}{N}.$ (3.17)
The second one
$\begin{split}\left\lvert\operatorname{\mathbf{E}}\left\{\frac{(A_{j}^{(0)}-A_{j})^{2}}{\bigl(A_{j}^{(0)}\bigr)^{2}A_{j}}\right\}\right\rvert&\leq\operatorname{\mathbf{E}}\left\{\left\lvert\frac{(A_{j}^{(0)}-A_{j})^{2}}{\bigl(A_{j}^{(0)}\bigr)^{2}A_{j}}\right\rvert\right\}\leq\frac{\operatorname{\mathbf{E}}\left\{\left\lvert A_{j}^{(0)}-A_{j}\right\rvert^{2}\right\}}{\left\lvert\Im z\right\rvert^{3}}\\ &\leq C\biggl(\frac{1}{N}\operatorname{\mathbf{E}}\left\{\left\lvert R_{jj}\right\rvert^{2}\right\}+\sum\limits_{i=1}^{4}\operatorname{\mathbf{E}}\left\{\left\lvert r_{i}\right\rvert^{2}\right\}\biggr).\end{split}$ (3.18)
Using Lemma 3.2, one gets
$\begin{split}\left\lvert\operatorname{\mathbf{E}}\left\{\frac{(A_{j}^{(0)}-A_{j})^{2}}{\bigl(A_{j}^{(0)}\bigr)^{2}A_{j}}\right\}\right\rvert&\leq\frac{C}{N}.\end{split}$ (3.19)
Therefore, $\left\lvert\operatorname{\mathbf{E}}\left\{g_{\mu}(z)\right\}-\operatorname{\mathbf{E}}\left\{g_{\nu}(z+\sigma^{2}\operatorname{\mathbf{E}}\left\{g_{\mu}(z)\right\})\right\}\right\rvert\leq\frac{C}{N}$. ∎
Proposition 3.3.
Let $W$ be defined in (1.2). Let $R$ be a square real symmetric matrix whose entries are i.i.d. random variables with zero mean and variance $\sigma^{2}$, and let the ESD $\nu$ of $S$ converge weakly to $\nu_{0}$ as $N\to\infty$. Then the Stieltjes transform $g_{\mu}(z)$ satisfies the following equation
$\operatorname{\mathbf{E}}g_{\mu}(z)-\operatorname{\mathbf{E}}g_{\nu}(z+\sigma^{2}\operatorname{\mathbf{E}}g_{\mu}(z))=O(N^{-1}),\quad N\to\infty.$ (3.20)
Proof.
Introduce a parametrized family of matrices $W(t)$ in the following way:
$W(t)=\dfrac{1}{\sqrt{N}}R(t)+S$ (3.21)
$R(t)=\sqrt{t}R+\sqrt{1-t}H.$ (3.22)
where entries of $R$ are i.i.d. random variables, with mean zero and variance $\sigma^{2}$, and $H$ is a Gaussian matrix (GOE), whose entries have mean zero and variance $\sigma^{2}$.
Note that for $t=0$, $R(0)=H$ with Gaussian entries, and we proved (3.20) for $t=0$, in Proposition 3.1. Here, we need to show that the same equation holds for $W(1)$, when the random matrix $R(1)=R$ is a Wigner matrix.
Introduce $G(t,z)$, the resolvent of $W(t)$
$G(t,z)=(W(t)-z)^{-1}$ (3.23)
and the Stieltjes transform $g_{\mu}(t,z)$
$g_{\mu}(t,z)=\dfrac{1}{N}\operatorname{\mathbf{E}}\{\operatorname{Tr}G(t,z)\}=\dfrac{1}{N}\sum\operatorname{\mathbf{E}}\left\{G_{ii}(t,z)\right\}.$ (3.24)
We prove the following lemma, first
Lemma 3.4.
$\left|\dfrac{dg_{\mu}(t,z)}{dt}\right|=O\left(\dfrac{1}{N}\right)$ uniformly in $t\in[0,1]$.
The proof of the lemma is given in Appendix B. Now, Lemma 3.4 implies
$\int_{0}^{1}\dfrac{dg_{\mu}(t,z)}{dt}dt=g_{\mu}(1,z)-g_{\mu}(0,z)=O\left(\dfrac{1}{N}\right)$ (3.25)
Thus,
$g_{\mu}(1,z)=g_{\mu}(0,z)+O\left(\dfrac{1}{N}\right)$ (3.26)
where $g_{\mu}(1,z)=g_{\mu}(z)$ is an object of interest in (3.20) and for $g_{\mu}(0,z)$ the same statement was proven in Prop.3.1. As the (3.2) and (3.20) both hold up to an order $O\left(\dfrac{1}{N}\right)$, and $g_{\mu}(1,z)$ and $g_{\mu}(0,z)$ differ from each other by the term of the same order, we conclude that (3.20) holds for $g_{\mu}(1,z)$. ∎
4 The limiting distribution of eigenvalues outside the bulk
In this section, we derive the limiting ESD of outliers. We will need the following lemma:
Lemma 4.1.
The functions $\Phi$ and $\omega_{\mu_{0}}$ defined in (1.13) and (1.8), respectively, satisfy the following: $\Phi(\omega_{\mu_{0}}(z))=z$ for any $z\in\mathbb{C}^{+}\setminus\operatorname{supp}{\mu_{0}}$.
Proof.
Use $\omega_{\mu_{0}}(z)$ (defined in (1.8)) and $\Phi(z)$ (defined in (1.13)) to compute $\Phi(\omega_{\mu_{0}}(z))$
$\Phi(\omega_{\mu_{0}}(z))=\omega_{\mu_{0}}(z)-\sigma^{2}g_{\nu_{0}}(\omega_{\mu_{0}}(z))$ (4.1)
Use the right hand side of (1.8) for measure $\tau=\mu_{0}$ to get
$\Phi(\omega_{\mu_{0}}(z))=z+\sigma^{2}g_{\mu_{0}}(z)-\sigma^{2}g_{\nu_{0}}(\omega_{\mu_{0}}(z))$ (4.2)
Note that $g_{\mu_{0}}(z)=g_{\nu_{0}}(\omega_{\mu_{0}}(z))$ by (1.7). Thus, we show that for any $z\in\mathbb{C}^{+}\setminus\operatorname{supp}{\mu_{0}}$, $\Phi(\omega_{\mu_{0}}(z))=z$. ∎
Set
$\tilde{\nu}_{1}$ $=\frac{N}{r}(\nu-\nu_{0}),$ (4.3)
$\tilde{\mu}_{1}$ $=\frac{N}{r}(\mu-\mu_{0}).$ (4.4)
Lemma 4.2.
Assume that the signed measure $\tilde{\nu}_{1}$ weakly converges to a signed measure $\nu_{1}$. Then the signed measure $\tilde{\mu}_{1}$ also converges weakly in probability to a signed measure $\mu_{1}$ whose Stieltjes transform $g_{\mu_{1}}(z)$ satisfies
$g_{\mu_{1}}(z)=g_{\nu_{1}}(\omega_{\mu_{0}}(z))\omega^{\prime}_{\mu_{0}}(z)$ (4.5)
for any $z\in\mathbb{C}^{+}$.
Proof.
Note that (3.2) can be presented in the form
$f_{\mu}(z)=f_{\nu}(\operatorname{\mathbf{E}}\omega_{\mu}(z))+O(N^{-1}),$ (4.6)
where
$f_{\tau}(z):=\operatorname{\mathbf{E}}\left\{g_{\tau}(z)\right\}.$ (4.7)
Expand both sides of (4.6). Taking into account that
$\omega_{\mu}(z)=\omega_{\mu_{0}}(z)+\frac{r}{N}\sigma^{2}g_{\tilde{\mu}_{1}}(z),$
we get
$\begin{split}f_{\mu_{0}}(z)+\tfrac{r}{N}f_{\tilde{\mu}_{1}}(z)={}&f_{\nu_{0}}(\operatorname{\mathbf{E}}\omega_{\mu_{0}}(z))\\ &+\tfrac{r}{N}\Bigl\{f^{\prime}_{\nu_{0}}(\operatorname{\mathbf{E}}\omega_{\mu_{0}}(z))\sigma^{2}f_{\tilde{\mu}_{1}}(z)\\ &+f_{\tilde{\nu}_{1}}(\operatorname{\mathbf{E}}\omega_{\mu_{0}}(z))\Bigr\}+o\left(\tfrac{r}{N}\right).\end{split}$ (4.8)
Since $f_{\mu_{0}}(z)=f_{\nu_{0}}(\operatorname{\mathbf{E}}\omega_{\mu_{0}}(z))$, we have
$f_{\tilde{\mu}_{1}}(z)=f^{\prime}_{\nu_{0}}(\operatorname{\mathbf{E}}\omega_{\mu_{0}}(z))\sigma^{2}f_{\tilde{\mu}_{1}}(z)+f_{\tilde{\nu}_{1}}(\operatorname{\mathbf{E}}\omega_{\mu_{0}}(z))+o(1).$ (4.9)
Note that $\mu_{0}$ and $\nu_{0}$ are non-random measures, thus we can remove expected values from (4.9) and get
$f_{\tilde{\mu}_{1}}(z)=g^{\prime}_{\nu_{0}}(\omega_{\mu_{0}}(z))\sigma^{2}f_{\tilde{\mu}_{1}}(z)+f_{\tilde{\nu}_{1}}(\omega_{\mu_{0}}(z))+o(1).$ (4.10)
Solving (4.10) for $f_{\tilde{\mu}_{1}}(z)$, we obtain
$f_{\tilde{\mu}_{1}}(z)=f_{\tilde{\nu}_{1}}(\omega_{\mu_{0}}(z))\dfrac{1}{1-\sigma^{2}g^{\prime}_{\nu_{0}}(\omega_{\mu_{0}}(z))}+o(1).$ (4.11)
Now let us prove that the measure $\tilde{\mu}_{1}$ has a weak limit as $N\to\infty$. It is sufficient to check that
  1. (i)
    $\int_{M}^{\infty}d\tilde{\mu}_{1}$ uniformly in $N$, in probability goes to zero as $M\to\infty$;
  2. (ii)
    $\operatorname{Var}\{g_{\tilde{\mu}_{1}}(z)\}\to 0$ as $N\to\infty$;
  3. (iii)
    $f_{\tilde{\mu}_{1}}(z)$ converges uniformly on compacts as $N\to\infty$.
(iii) follows from Proposition 3.1. (ii) follows from the Poincaré inequality. Indeed, consider $g_{\tilde{\mu}_{1}}(z)$ as a function of $R$
$\Psi(R)=\Psi_{z}(R):=g_{\tilde{\mu}_{1}}(z)=\frac{1}{r}\operatorname{Tr}\left(N^{-1/2}R+S-z\right)^{-1}-\frac{N}{r}g_{\mu_{0}}(z).$ (4.12)
The Poincaré inequality implies
$\operatorname{Var}\{g_{\tilde{\mu}_{1}}(z)\}=\operatorname{Var}\{\Psi(R)\}\leq\operatorname{\mathbf{E}}\left\{\sigma^{2}\left\lvert\nabla\Psi(R)\right\rvert^{2}\right\}.$ (4.13)
Next, let us calculate the derivatives of $\Psi$. Differentiating the identity (Wz)G(z)=I(W-z)G(z)=I, where $G(z)$ is defined in (3.5), one gets
$\frac{\partial}{\partial R_{jk}}G(z)$ $=-N^{-1/2}G(z)(E_{jk}+E_{kj})G(z),\quad j<k,$ (4.14)
$\frac{\partial}{\partial R_{jj}}G(z)$ $=-N^{-1/2}G(z)E_{jj}G(z)$ (4.15)
with $E_{jk}$ be coordinate matrices, i.e. all their entries are zeros except one unit in the $j$th row and $k$th column. Taking trace of both sides and dividing by $r$, we have
$\frac{\partial}{\partial R_{jk}}\Psi(R)$ $=-\frac{1}{r\sqrt{N}}((G^{2})_{jk}+(G^{2})_{kj}),\quad j<k,$ (4.16)
$\frac{\partial}{\partial R_{jj}}\Psi(R)$ $=-\frac{1}{r\sqrt{N}}(G^{2})_{jj}.$ (4.17)
Therefore,
$\operatorname{\mathbf{E}}\left\{\sigma^{2}\left\lvert\nabla\Psi(R)\right\rvert^{2}\right\}\leq\frac{2\sigma^{2}}{r^{2}N}\operatorname{Tr}(G(z))^{4}\leq\frac{2\sigma^{2}}{r^{2}}\left\lVert(G(z))^{4}\right\rVert\leq\frac{2\sigma^{2}}{r^{2}\left\lvert\Im z\right\rvert^{4}}.$ (4.18)
Since $r\to\infty$ as $N\to\infty$, the combination of (4.13) and (4.18) yields (ii).
To prove (i), consider
$v_{2}:=\int t^{2}d\tilde{\mu}_{1}(t)=\frac{N}{r}\left(\int t^{2}d\mu(t)-\int t^{2}d\mu_{0}(t)\right)=\frac{N}{r}\left(\operatorname{Tr}W^{2}-\int t^{2}d\mu_{0}(t)\right).$ (4.19)
The second moment of the measure $\mu_{0}$ can be recovered from the equation on its Stieltjes transform (1.7)
$\int t^{2}d\mu_{0}(t)=\sigma^{2}+\int t^{2}d\nu_{0}(t).$ (4.20)
Then,
$\operatorname{Tr}W^{2}=\operatorname{Tr}\left(\frac{1}{\sqrt{N}}R+S\right)^{2}=\frac{1}{N}\operatorname{Tr}R^{2}+\frac{2}{\sqrt{N}}\operatorname{Tr}RS+\operatorname{Tr}S^{2}.$ (4.21)
Substituting (4.20) and (4.21) into (4.19), one obtains
$\begin{split}v_{2}&=\frac{N}{r}\left[\frac{1}{N}\operatorname{Tr}R^{2}-\sigma^{2}\right]+\frac{2\sqrt{N}}{r}\operatorname{Tr}RS+\frac{N}{r}\left[\operatorname{Tr}S^{2}-\int t^{2}d\nu_{0}(t)\right]\\ &=:\Sigma_{1}+\Sigma_{2}+\Sigma_{3}.\end{split}$ (4.22)
$\Sigma_{3}$ converges to $\int t^{2}d\nu_{1}$ by Assumption 2 and
$P(\left\lvert\Sigma_{i}\right\rvert>\varepsilon)\to 0,\quad i=1,2$ (4.23)
by the weak law of large numbers. Hence, as $N\to\infty$, $v_{2}$ converges to $\int t^{2}d\nu_{1}$. It implies (i).
Now we can take the limit in (4.11). We get
$g_{\mu_{1}}(z)=g_{\nu_{1}}(\omega_{\mu_{0}}(z))\dfrac{1}{1-\sigma^{2}g^{\prime}_{\nu_{0}}(\omega_{\mu_{0}}(z))}.$ (4.24)
Since $\Phi(\omega_{\mu_{0}}(z))=z$ for any $z\in\mathbb{C}^{+}\setminus\operatorname{supp}{\mu_{0}}$, where $\Phi$ is defined in (1.13), it is sufficient to verify that the denominator of (4.24) is equal to $\Phi^{\prime}(\omega_{\mu_{0}}(z))$, i.e.
$\Phi^{\prime}(\omega_{\mu_{0}}(z))=1-\sigma^{2}g^{\prime}_{\nu_{0}}(\omega_{\mu_{0}}(z)),$ (4.25)
Taking derivative in (1.13) we obtain
$\Phi^{\prime}(z)=1-\sigma^{2}g^{\prime}_{\nu_{0}}(z)$ (4.26)
Substituting $z\to\omega_{\mu_{0}}(z)$ into the equation above, we get (4.25).
By Lemma 4.1, $\Phi(\omega_{\mu_{0}}(z))=z$ for any $z\in\mathbb{C}^{+}\setminus\operatorname{supp}{\mu_{0}}$. Hence,
$\Phi^{\prime}(\omega_{\mu_{0}}(z))=\dfrac{1}{\omega^{\prime}_{\mu_{0}}(z)}.$ (4.27)
Use this fact and (4.24) to conclude that
$g_{\mu_{1}}(z)=g_{\nu_{1}}(\omega_{\mu_{0}}(z))\omega^{\prime}_{\mu_{0}}(z).$ (4.28)
Next, we present the proof of Theorem 2.1.
Proof of Theorem 2.1.
Let $z=x+iy$, $x\notin\operatorname{supp}\mu_{0}$.
As we have shown before in Lemma 4.2, the signed measure $\tilde{\mu}_{1}=\frac{N}{r}(\mu-\mu_{0})$ converges weakly in probability to the signed measure $\mu_{1}$ and its Stieltjes transform satisfies
$g_{\mu_{1}}(z)=g_{\nu_{1}}(\omega_{\mu_{0}}(z))\omega^{\prime}_{\mu_{0}}(z).$ (4.29)
Then,
$\begin{split}\rho_{\mu_{1}}(x)&=\lim_{y\to+0}\dfrac{1}{\pi}\Im g_{\mu_{1}}(x+iy)=\lim_{y\to+0}\dfrac{1}{\pi}\Im\left(g_{\nu_{1}}(\omega_{\mu_{0}}(x+iy))\omega^{\prime}_{\mu_{0}}(x+iy)\right)\\ &=\lim_{y\to+0}\dfrac{1}{\pi}\left(\Im g_{\nu_{1}}(\omega_{\mu_{0}}(x+iy))\Re\omega^{\prime}_{\mu_{0}}(x+iy)+\Re g_{\nu_{1}}(\omega_{\mu_{0}}(x+iy))\Im\omega^{\prime}_{\mu_{0}}(x+iy)\right)\end{split}$ (4.30)
The definition (1.8) implies that $\omega_{\mu_{0}}(z)$ is a holomorphic function for any $z\in\mathbb{C}\setminus\operatorname{supp}{\mu_{0}}$, thus it is continuous, and it’s derivative $\omega^{\prime}_{\mu_{0}}(z)$ is also continuous on the same set. Thus,
$\lim_{y\to+0}\Re\omega_{\mu_{0}}^{\prime}(x+iy)$ $=\Re\omega_{\mu_{0}}^{\prime}(x),$ (4.31)
$\lim_{y\to+0}\omega_{\mu_{0}}(x+iy)$ $=\omega_{\mu_{0}}(x)$ (4.32)
i.e., $\lim_{y\to+0}\Im\omega_{\mu_{0}}(x+iy)=0$ and $\lim_{y\to+0}\Re\omega_{\mu_{0}}(x+iy)=\omega_{\mu_{0}}(x)$. Changing variables $t=\Im\omega_{\mu_{0}}(x+iy)$, we get
$\lim_{y\to+0}\frac{1}{\pi}g_{\nu_{1}}(\omega_{\mu_{0}}(x+iy))=\lim_{t\to 0}\frac{1}{\pi}g_{\nu_{1}}(\omega_{\mu_{0}}(x)+o(t)+it).$ (4.33)
Let us check that $\omega_{\mu_{0}}(x+iy)$ has a positive imaginary part. From the definition (1.8) it follows that
$\Im\omega_{\mu_{0}}(x+iy)=\Im((x+iy)+\sigma^{2}g_{\mu_{0}}(x+iy))=y+\sigma^{2}\Im g_{\mu_{0}}(x+iy),$ (4.34)
By properties of the Stieltjes transform, $y\Im g_{\mu_{0}}(x+iy)>0$. Therefore, $\Im g_{\mu_{0}}(x+iy)>0$ for $y>0$. Also, for $x\notin\operatorname{supp}\mu_{0}$, there exists $\delta>0$ such that $(x-\delta,x+\delta)\cap\operatorname{supp}\mu_{0}=\emptyset$. We have for $0<y<\delta^{2}$
$\begin{split}\left\lvert\Im g_{\mu_{0}}(x+iy)\right\rvert=\left\lvert\int_{\left\lvert t-x\right\rvert\geq\delta}\frac{yd\mu_{0}(t)}{(t-x)^{2}+y^{2}}\right\rvert\leq C\int_{\left\lvert t-x\right\rvert\geq\delta}\frac{ydt}{(t-x)^{2}+y^{2}}\\ =C\int_{\left\lvert\tilde{t}\right\rvert\geq\delta/y}\frac{d\tilde{t}}{\tilde{t}^{2}+1}\leq C\frac{y}{\delta}<C\sqrt{y}.\ \end{split}$ (4.35)
Hence, $\Im\omega_{\mu_{0}}(x+iy)>0$, as a sum of two positive numbers, and
$\Im\omega_{\mu_{0}}(x+iy)\leq C\sqrt{y}\quad\text{as }y\to+0.$ (4.36)
Then, (4.33) transforms to
$\lim_{y\to+0}\frac{1}{\pi}\Im g_{\nu_{1}}(\omega_{\mu_{0}}(x+iy))=\lim_{t\to+0}\frac{1}{\pi}\Im g_{\nu_{1}}(\omega_{\mu_{0}}(x)+o(t)+it)=\rho_{\nu_{1}}(\omega_{\mu_{0}}(x)).$ (4.37)
Let us consider the second term in (4.30), $\Re g_{\nu_{1}}(\omega_{\mu_{0}}(x+iy))\Im\omega^{\prime}_{\mu_{0}}(x+iy)$. For $w=x_{1}+iy_{1}$ one gets
$\Re g_{\nu_{1}}(w)=\int\frac{(t-x_{1})d\nu_{1}(t)}{(t-x_{1})^{2}+y_{1}^{2}}$ (4.38)
and
$\begin{split}\left\lvert\Re g_{\nu_{1}}(w)\right\rvert\leq\left\lvert\int_{\left\lvert t-x_{1}\right\rvert\leq y_{1}}+\int_{\left\lvert t-x_{1}\right\rvert>y_{1}}\frac{(t-x_{1})d\nu_{1}(t)}{(t-x_{1})^{2}+y_{1}^{2}}\right\rvert\\ \leq\int_{\left\lvert t-x_{1}\right\rvert\leq y_{1}}\frac{d\left\lvert\nu_{1}\right\rvert(t)}{y_{1}}+\int_{\left\lvert t-x_{1}\right\rvert>y_{1}}\frac{d\left\lvert\nu_{1}\right\rvert(t)}{\left\lvert t-x_{1}\right\rvert}\leq C-C\log y_{1}.\end{split}$ (4.39)
Thus, taking into account (4.36), one obtains, as $y\to+0$,
$\left\lvert\Re g_{\nu_{1}}(\omega_{\mu_{0}}(x+iy))\right\rvert\leq C(1-\log\Im\omega_{\mu_{0}}(x+iy))\leq C(1-\log y).$ (4.40)
Next,
$\Im\omega^{\prime}_{\mu_{0}}(x+iy)=\sigma^{2}\Im g_{\mu_{0}}^{\prime}(x+iy)=2\sigma^{2}\int\frac{(t-x)yd\mu_{0}(t)}{((t-x)^{2}+y^{2})^{2}}.$ (4.41)
It implies
$\lim\limits_{y\to+0}\frac{\Im\omega^{\prime}_{\mu_{0}}(x+iy)}{y}=\sigma^{2}\int\frac{d\mu_{0}(t)}{(t-x)^{3}}.$ (4.42)
Hence, $\Im\omega^{\prime}_{\mu_{0}}(x+iy)\leq Cy$ and
$\left\lvert\Re g_{\nu_{1}}(\omega_{\mu_{0}}(x+iy))\Im\omega^{\prime}_{\mu_{0}}(x+iy)\right\rvert\leq C(1-\log y)y\to 0\quad\text{as }y\to+0.$ (4.43)
A combination of (4.30), (4.37), and (4.43) gives us
$\rho_{\mu_{1}}(x)=\rho_{\nu_{1}}(\omega_{\mu_{0}}(x))\omega_{\mu_{0}}^{\prime}(x).$ (4.44)
Finally,
$\mu_{1}(\Delta)=\int_{\Delta}\rho_{\mu_{1}}(x)dx=\int_{\Delta}\rho_{\nu_{1}}(\omega_{\mu_{0}}(x))\omega_{\mu_{0}}^{\prime}(x)dx=\int_{\omega_{\mu_{0}}(\Delta)}\rho_{\nu_{1}}(x_{1})dx_{1}=\nu_{1}(\omega_{\mu_{0}}(\Delta)).$ (4.45)
5 Proof of Theorem 2.2
In this section, we present the proof of Theorem 2.2.
According to Theorem 2.1, $\operatorname{supp}\mu_{1}^{+}=[a,b]$. Pick an arbitrary positive integer $k$. Let the points $\zeta_{1}\leq\dotsb\leq\zeta_{k-1}$ split the segment $[a,b]$ into $k$ equal parts, i.e.
$\zeta_{i}=a+\frac{i}{k}(b-a),\quad 0<i<k.$ (5.1)
The continuity of $\Phi(x)$, Lemma 4.1, and (2.4) imply $\lim_{N\to\infty}\max_{1\leq j\leq r}\operatorname{dist}(\lambda_{j}(W),[a,b])=0$ and $\lambda_{1}(W)\leq c$ for some $c$. Fix an arbitrary $\varepsilon_{1},\varepsilon_{2}>0$ and consider functions
$h_{i}(x)=\begin{cases}0,&x<\zeta_{i}\text{ or }x>c+\varepsilon_{1}\\ \frac{x-\zeta_{i}}{\varepsilon_{1}},&\zeta_{i}\leq x<\zeta_{i}+\varepsilon_{1}\\ 1,&\zeta_{i}+\varepsilon_{1}\leq x<c\\ \frac{c+\varepsilon_{1}-x}{\varepsilon_{1}},&c\leq x\leq c+\varepsilon_{1}\end{cases}$ (5.2)
[…]
Take any positive $\varepsilon_{3}$ such that
$\varepsilon_{3}<\frac{1}{2}\min_{i}\int(h_{i}(x)-h_{i+1}(x))d\mu_{1}(x).$ (5.3)
Assumption 3 and Theorem 2.1 yield that there exists $N_{0}$ such that for any $N>N_{0}$
$\sum_{i}\left\lvert\tfrac{N}{r}\int h_{i}(\Phi(x))d(\nu(x)-\nu_{0}(x))-\int h_{i}(\Phi(x))d\nu_{1}(x)\right\rvert<\varepsilon_{3},$ (5.4)
$P\left(\sum_{i}\left\lvert\tfrac{N}{r}\int h_{i}(x)d\mu(x)-\int h_{i}(x)d\mu_{1}(x)\right\rvert>\varepsilon_{3}\right)<\varepsilon_{2}.$ (5.5)
Fix some $N>N_{0}$ and let $\lambda_{j(N)}(W)\in[\zeta_{m},\zeta_{m+1}]$. Using (5.4), we get
$\begin{split}\#\{{i\mid\lambda_{i}(S)\geq\omega_{\mu_{0}}(\zeta_{m+3})}\}&=N\int 1_{\omega_{\mu_{0}}(\zeta_{m+3})\leq x\leq c}(x)d\nu(x)\leq N\int h_{m+2}(\Phi(x))d\nu(x)\\ &<r\int h_{m+2}(\Phi(x))d\nu_{1}(x)+r\varepsilon_{3}.\end{split}$ (5.6)
By Theorem 2.1 and by the choice of $\varepsilon_{3}$ (5.3),
$\int h_{m+2}(\Phi(x))d\nu_{1}(x)=\int h_{m+2}(x)d\mu_{1}(x)<\int h_{m+1}(x)d\mu_{1}(x)-2\varepsilon_{3}$ (5.7)
(5.5) implies
$\begin{split}r\int h_{m+1}(x)d\mu_{1}(x)-r\varepsilon_{3}<N\int h_{m+1}(x)d\mu(x)\leq\#\{{i\mid\lambda_{i}(W)>\zeta_{m+1}}\}<j(N)\end{split}$ (5.8)
with probability greater or equal to $1-\varepsilon_{2}$. Hence, combining (5.6)–(5.8) one obtains $\lambda_{j(N)}(S)<\omega_{\mu_{0}}(\zeta_{m+3})$. Similarly,
$\begin{split}\#\{{i\mid\lambda_{i}(S)\geq\omega_{\mu_{0}}(\zeta_{m-2})}\}&\geq N\int h_{m-2}(\Phi(x))d\nu(x)>r\int h_{m-2}(\Phi(x))d\nu_{1}(x)-r\varepsilon_{3}\\ &>r\int h_{m-1}(x)d\mu_{1}(x)+r\varepsilon_{3}>N\int h_{m-1}(x)d\mu(x)\\ &\geq\#\{{i\mid\lambda_{i}(W)>\zeta_{m}}\}\geq j(N)\end{split}$ (5.9)
and $\omega_{\mu_{0}}(\zeta_{m-2})<\lambda_{j(N)}(S)$. Thus, $\zeta_{m-2}<\Phi(\lambda_{j(N)}(S))<\zeta_{m+3}$ which implies
$P\left[\left\lvert\lambda_{j(N)}(W)-\Phi(\lambda_{j(N)}(S))\right\rvert<\frac{3}{k}(b-a)\right]>1-\varepsilon_{2}.$ (5.10)
Since $k$ and $\varepsilon_{2}$ are arbitrary, the last equation yields the convergence (2.5) in probability.
Appendix A Numerical simulations
For the numerical simulations below, the following DNN was considered:
Networks were trained for 25 epochs on the Fashion MNIST dataset. Each network has 3 weight layers of the following sizes: $784\times N$, $N\times N$, $N\times 10$, where the matrix size $N$ changes from 200 to 3000, with step 200. For every $N$ the network achieves test accuracy 88-89%.  [^1] The code can be found at GitHub by the following link.
Figure 2: Numerical simulations of a DNN with 3 layers. Two figures correspond to two different realizations with the same architecture.
These numerical simulations demonstrate the dependence of the number of eigenvalues of the ‘‘signal’’ matrix on the size of the middle layer matrix.
Appendix B Proof of auxiliary lemmas
Proof of Lemma 3.2.
Let $\operatorname{\mathbf{E}}_{i}$ be an expectation with respect to $R_{ij}$, $j=1,\dotsc,N$. We have
$\operatorname{\mathbf{E}}_{i}\left\{r_{1}\overline{r_{1}}\right\}=\dfrac{1}{N^{2}}\sum_{j_{1}\neq k_{1};\ j_{1},k_{1}\neq i}^{N}\sum_{j_{2}\neq k_{2};\ j_{2},k_{2}\neq i}^{N}G_{j_{1}k_{1}}^{(i)}\overline{G}_{j_{2}k_{2}}^{(i)}\operatorname{\mathbf{E}}_{i}\left\{R_{ij_{1}}\overline{R}_{ik_{1}}R_{ij_{2}}\overline{R}_{ik_{2}}\right\}.$ (B.1)
If $j_{1}\neq j_{2}$ and $j_{1}\neq k_{2}$, then $R_{ij_{1}}$ and $\overline{R}_{ik_{1}}R_{ij_{2}}\overline{R}_{ik_{2}}$ are independent, which implies
$\operatorname{\mathbf{E}}_{i}\left\{R_{ij_{1}}\overline{R}_{ik_{1}}R_{ij_{2}}\overline{R}_{ik_{2}}\right\}=\operatorname{\mathbf{E}}_{i}\left\{R_{ij_{1}}\right\}\operatorname{\mathbf{E}}_{i}\left\{\overline{R}_{ik_{1}}R_{ij_{2}}\overline{R}_{ik_{2}}\right\}=0.$ (B.2)
Similarly, the expectation is also zero if $k_{1}\neq j_{2}$ and $k_{1}\neq k_{2}$. Hence,
$\begin{split}\operatorname{\mathbf{E}}_{i}\left\{r_{1}\overline{r_{1}}\right\}&=\dfrac{1}{N^{2}}\sum_{\begin{subarray}{c}j_{1}\neq k_{1}\\ j_{1},k_{1}\neq i\end{subarray}}G_{j_{1}k_{1}}^{(i)}\overline{G}_{k_{1}j_{1}}^{(i)}\operatorname{\mathbf{E}}_{i}\left\{\left\lvert R_{ij_{1}}R_{ij_{2}}\right\rvert^{2}\right\}\\ &\quad{}+\dfrac{1}{N^{2}}\sum_{\begin{subarray}{c}j_{1}\neq k_{1}\\ j_{1},k_{1}\neq i\end{subarray}}^{N}\left\lvert G_{j_{1}k_{1}}^{(i)}\right\rvert^{2}\operatorname{\mathbf{E}}_{i}\left\{R_{ij_{1}}^{2}\overline{R}_{ik_{1}}^{2}\right\}\\ &\leq\frac{2}{N^{2}}\operatorname{Tr}G^{(i)}(G^{(i)})^{*}\leq\frac{2}{N}\left\lVert G^{(i)}\right\rVert^{2}\leq\frac{2}{N\left\lvert\Im z\right\rvert^{2}}.\end{split}$ (B.3)
So, $\operatorname{\mathbf{E}}\left\{\left\lvert r_{1}\right\rvert^{2}\right\}\leq\frac{C}{N}$. Let us proceed to derive an estimate on $r_{2}$. We get
$\begin{split}\operatorname{\mathbf{E}}_{i}\left\{r_{2}\overline{r_{2}}\right\}&=\dfrac{1}{N^{2}}\sum_{j_{1}\neq i}\sum_{j_{2}\neq i}G_{j_{1}j_{1}}^{(i)}\overline{G}_{j_{2}j_{2}}^{(i)}\operatorname{\mathbf{E}}_{i}\left\{(\left\lvert R_{ij_{1}}\right\rvert^{2}-\sigma^{2})(\left\lvert R_{ij_{2}}\right\rvert^{2}-\sigma^{2})\right\}\\ &=\dfrac{1}{N^{2}}\sum_{j\neq i}\left\lvert G_{jj}^{(i)}\right\rvert^{2}\operatorname{\mathbf{E}}_{i}\left\{(\left\lvert R_{ij}\right\rvert^{2}-\sigma^{2})^{2}\right\}\leq\frac{c_{4}}{N^{2}}\operatorname{Tr}G^{(i)}(G^{(i)})^{*}\leq\frac{c_{4}}{n\left\lvert\Im z\right\rvert^{2}},\end{split}$ (B.4)
because $\operatorname{\mathbf{E}}_{i}\left\{(\left\lvert R_{ij}\right\rvert^{2}-\sigma^{2})^{2}\right\}\leq c_{4}$. Further, $r_{3}$ can be written as follows
$\operatorname{\mathbf{E}}\left\{r_{3}\overline{r_{3}}\right\}=\operatorname{Var}\left\{\frac{1}{N}\operatorname{Tr}G^{(i)}\right\}\leq\frac{C}{N},$ (B.5)
where the last inequality is given in (18.2.13) of Theorem 18.2.3 in [8]. Since
$\left\lvert\operatorname{Tr}G^{(i)}-\operatorname{Tr}G\right\rvert\leq\frac{1}{\left\lvert\Im z\right\rvert}$, (B.6)
one has $r_{4}\leq\frac{C}{N}$. ∎
Proof of Lemma 3.4.
Take a derivative of $g(t,z)$ with respect to $t$:
$\dfrac{\partial}{\partial t}g_{\mu}(t,z)=\dfrac{1}{N}\operatorname{\mathbf{E}}\{\operatorname{Tr}\dfrac{\partial}{\partial t}G(t,z)\}$ (B.7)
$\dfrac{\partial}{\partial t}G(t,z)=-G\cdot\dfrac{\partial}{\partial t}W(t,z)\cdot G$ (B.8)
Using (3.21) and the fact that $S$ do not depend on $t$, we get
$\dfrac{\partial}{\partial t}W(t,z)=\frac{1}{2\sqrt{t}\sqrt{N}}R-\frac{1}{2\sqrt{1-t}\sqrt{N}}H$ (B.9)
Thus,
$\dfrac{\partial}{\partial t}g_{\mu}(t,z)=-\dfrac{1}{2N}\left(\dfrac{1}{\sqrt{t}\sqrt{N}}\operatorname{\mathbf{E}}\{\operatorname{Tr}(GRG)\}-\dfrac{1}{\sqrt{1-t}\sqrt{N}}\operatorname{\mathbf{E}}\{\operatorname{Tr}(GHG)\}\right)$ (B.10)
Rewrite (B.10) in terms of the entries of $R$, $H$, and $G$:
$\operatorname{\mathbf{E}}\{\operatorname{Tr}(GRG)\}=\operatorname{\mathbf{E}}\{\operatorname{Tr}(RG^{2})\}=\operatorname{\mathbf{E}}\{\sum_{i,j=1}^{N}(R)_{ij}(G^{2})_{ji}\}=\sum_{i,j=1}^{N}\operatorname{\mathbf{E}}\{(R)_{ij}(G^{2})_{ji}\}$ (B.11)
and following the same steps, get a similar representation for the Gaussian term, then (B.10) becomes
$\dfrac{\partial}{\partial t}g_{\mu}(t,z)=-\dfrac{1}{2N}\left(\dfrac{1}{\sqrt{t}\sqrt{N}}\sum_{i,j=1}^{N}\operatorname{\mathbf{E}}\{(R)_{ij}(G^{2})_{ji}\}-\dfrac{1}{\sqrt{1-t}\sqrt{N}}\sum_{i,j=1}^{N}\operatorname{\mathbf{E}}\{(H)_{ij}(G^{2})_{ji}\}\right)$ (B.12)
For the Wigner term above, use Proposition 18.1.4 from [8] for $p=2$, random variable $(R)_{ij}$, and a function $(G^{2})_{ji}$ to get the following decompositions for any $1\leq i,j\leq N$:
$\operatorname{\mathbf{E}}\{(R)_{ij}(G^{2})_{ji}\}=\kappa_{1}^{R}\operatorname{\mathbf{E}}\{(G^{2})_{ji}\}+\kappa_{2}^{R}\operatorname{\mathbf{E}}\{\dfrac{\partial(G^{2})_{ji}}{\partial(R)_{ij}}\}+\dfrac{\kappa_{3}^{R}}{2}\operatorname{\mathbf{E}}\{\dfrac{\partial^{2}(G^{2})_{ji}}{\partial(R)_{ij}^{2}}\}+\varepsilon_{ij},$ (B.13)
where $\kappa_{i}^{R}$ are the cumulants of $R$, and $\varepsilon_{ij}$ is bounded in the following way:
$|\varepsilon_{ij}|\leq C_{2}\operatorname{\mathbf{E}}\{|(R)_{ij}^{4}|\}\sup_{R_{ij}\in\mathbb{R}}\left\lvert\dfrac{\partial^{3}(G^{2})_{ji}}{\partial(R)_{ij}^{3}}\right\rvert$, (B.14)
and $C_{2}$ – constant. For the Gaussian term, use Lemma 2.1.5 from [8]:
$\operatorname{\mathbf{E}}\{(H)_{ij}(G^{2})_{ji}\}=\operatorname{\mathbf{E}}\{(H)_{ij}^{2}\}\operatorname{\mathbf{E}}\{\dfrac{\partial(G^{2})_{ji}}{\partial(H)_{ij}}\},$ (B.15)
Compute the terms in (B.13) and (B.15) one by one, starting from the first derivatives.
In order to find $\operatorname{\mathbf{E}}\{\dfrac{\partial(G^{2})_{ji}}{\partial(R)_{ij}}\}$, use the defining property of the resolvent $G$
$G^{2}(W-zI)^{2}=1$ (B.16)
and differentiate it with respect to $(R)_{ij}$:
$\dfrac{\partial}{\partial(R)_{ij}}(G^{2}(W-zI)^{2})=0$ (B.17)
$\dfrac{\partial G^{2}}{\partial(R)_{ij}}(W-zI)^{2}+G^{2}\dfrac{\partial}{\partial(R)_{ij}}((W-zI)^{2})=0$ (B.18)
$\dfrac{\partial G^{2}}{\partial(R)_{ij}}(W-zI)^{2}+G^{2}\left(\dfrac{\partial W^{2}}{\partial(R)_{ij}}-2z\dfrac{\partial W}{\partial(R)_{ij}}\right)=0$ (B.19)
Here, $\dfrac{\partial W}{\partial(R)_{ij}}=\dfrac{\sqrt{t}}{\sqrt{N}}(E_{ij}+E_{ji})$, with $E_{ij}$ being coordinate matrices, i.e. all their entries are zeros except one unit in the $i$th row and the $j$th column, and
$\dfrac{\partial W^{2}}{\partial(R)_{ij}}=\dfrac{\partial W}{\partial(R)_{ij}}W+W\dfrac{\partial W}{\partial(R)_{ij}}=\dfrac{\sqrt{t}}{\sqrt{N}}\left((E_{ij}+E_{ji})W+W(E_{ij}+E_{ji})\right).$ (B.20)
Using derivatives of $W$ and $W^{2}$ in (B.19), we get
$\dfrac{\partial G^{2}}{\partial(R)_{ij}}(W-zI)^{2}+\dfrac{\sqrt{t}}{\sqrt{N}}G^{2}\left((E_{ij}+E_{ji})(W-zI)+(W-zI)(E_{ij}+E_{ji})\right)=0$ (B.21)
and multiplying both sides by $G^{2}$ from the right,
$\dfrac{\partial G^{2}}{\partial(R)_{ij}}=-\dfrac{\sqrt{t}}{\sqrt{N}}G^{2}(E_{ij}+E_{ji})G-\dfrac{\sqrt{t}}{\sqrt{N}}G(E_{ij}+E_{ji})G^{2}$ (B.22)
In order to find $\operatorname{\mathbf{E}}\{\dfrac{\partial(G^{2})_{ji}}{\partial(H)_{ij}}\}$, we repeat the same steps, and the final formula is almost the same as (B.22), with the only difference being the scalings as $\dfrac{\partial W}{\partial(H)_{ij}}=\dfrac{\sqrt{1-t}}{\sqrt{N}}E_{ij}$, thus
$\dfrac{\partial G^{2}}{\partial(H)_{ij}}=-\dfrac{\sqrt{1-t}}{\sqrt{N}}G^{2}(E_{ij}+E_{ji})G-\dfrac{\sqrt{1-t}}{\sqrt{N}}G(E_{ij}+E_{ji})G^{2}$ (B.23)
Now, find the second derivative $\operatorname{\mathbf{E}}\{\dfrac{\partial^{2}(G^{2})_{ji}}{\partial(R)^{2}_{ij}}\}$ by differentiating (B.18) one more time.
$\dfrac{\partial G^{2}}{\partial(R)_{ij}}=-G^{2}\dfrac{\partial}{\partial(R)_{ij}}((W-zI)^{2})G^{2}$ (B.24)
Differentiating it one more time, using the product rule:
$\dfrac{\partial^{2}G^{2}}{\partial(R)^{2}_{ij}}=-\dfrac{\partial G^{2}}{\partial(R)_{ij}}\cdot\dfrac{\partial}{\partial(R)_{ij}}((W-zI)^{2})G^{2}-G^{2}\dfrac{\partial^{2}}{\partial(R)^{2}_{ij}}((W-zI)^{2})G^{2}\\ -G^{2}\dfrac{\partial}{\partial(R)_{ij}}((W-zI)^{2})\dfrac{\partial G^{2}}{\partial(R)_{ij}}$ (B.25)
Substituting (B.24), we get
$\dfrac{\partial^{2}G^{2}}{\partial(R)^{2}_{ij}}=2G^{2}\dfrac{\partial}{\partial(R)_{ij}}((W-zI)^{2})G^{2}\dfrac{\partial}{\partial(R)_{ij}}((W-zI)^{2})G^{2}-G^{2}\dfrac{\partial^{2}}{\partial(R)^{2}_{ij}}((W-zI)^{2})G^{2}$ (B.26)
Now, compute the second derivative of $(W-zI)^{2}$. Since $\dfrac{\partial W}{\partial(R)_{ij}}=\dfrac{\sqrt{t}}{\sqrt{N}}(E_{ij}+E_{ji})$ is constant with respect to $R_{ij}$, $\dfrac{\partial^{2}W}{\partial(R)^{2}_{ij}}=0$, and
$\dfrac{\partial^{2}}{\partial(R)^{2}_{ij}}((W-zI)^{2})=\dfrac{\sqrt{t}}{\sqrt{N}}\dfrac{\partial}{\partial R_{ij}}\left((E_{ij}+E_{ji})(W-zI)+(W-zI)(E_{ij}+E_{ji})\right)=2\dfrac{t}{N}(E_{ij}+E_{ji})^{2}$ (B.27)
Also, from (B.21)
$\dfrac{\partial}{\partial(R)_{ij}}((W-zI)^{2})=\dfrac{\sqrt{t}}{\sqrt{N}}\left((E_{ij}+E_{ji})(W-zI)+(W-zI)(E_{ij}+E_{ji})\right)$ (B.28)
so
$\dfrac{\partial^{2}G^{2}}{\partial(R)^{2}_{ij}}=2\dfrac{t}{N}G^{2}\left((E_{ij}+E_{ji})(W-zI)+(W-zI)(E_{ij}+E_{ji})\right)G^{2}\left((E_{ij}+E_{ji})(W-zI)\right.\\ \left.+(W-zI)(E_{ij}+E_{ji})\right)G^{2}-2\dfrac{t}{N}G^{2}(E_{ij}+E_{ji})^{2}G^{2}$ (B.29)
Using that G(WzI)=1G(W-zI)=1 and canceling the identical terms:
$\dfrac{\partial^{2}G^{2}}{\partial(R)^{2}_{ij}}=2\dfrac{t}{N}\left(G^{2}(E_{ij}+E_{ji})G(E_{ij}+E_{ji})G\right.\\ \left.+G(E_{ij}+E_{ji})G^{2}(E_{ij}+E_{ji})G+G(E_{ij}+E_{ji})G(E_{ij}+E_{ji})G^{2}\right)$ (B.30)
Note that considering the $ji$-element of the matrix is the same as multiplying this matrix by $E_{ij}$ on the right and taking the trace. After that, use the cyclic property of the trace to unify the terms.
$\dfrac{\partial^{2}(G^{2})_{ji}}{\partial(R)^{2}_{ij}}=2\dfrac{t}{N}\operatorname{Tr}\left(G^{2}(E_{ij}+E_{ji})G(E_{ij}+E_{ji})GE_{ij}\right.\\ \left.+G^{2}(E_{ij}+E_{ji})GE_{ij}G(E_{ij}+E_{ji})+G^{2}E_{ij}G(E_{ij}+E_{ji})G(E_{ij}+E_{ji})\right)$ (B.31)
When using (B.13) and (B.15) for (B.12), note that $\kappa_{1}^{R}=0$ (as $R_{ij}$ have mean zero). Also, using (B.22), (B.23), and the fact that $\kappa_{2}^{R}=\operatorname{\mathbf{E}}\{(H)_{ij}^{2}\}$, we see that the $\kappa_{2}$-term is canceled with the Gaussian term in (B.12), thus
$\dfrac{\partial}{\partial t}g_{\mu}(t,z)=-\dfrac{1}{2\sqrt{t}N^{3/2}}\sum_{i,j=1}^{N}\left(\frac{\kappa_{3}^{R}}{2}\operatorname{\mathbf{E}}\{\dfrac{\partial^{2}(G^{2})_{ji}}{\partial(R)_{ij}^{2}}\}+\varepsilon_{ij}\right).$ (B.32)
Distribute the terms in (B.31) and take a sum over all $i,j$. We get four types of terms:
$\sum_{i,j=1}^{N}\operatorname{\mathbf{E}}\{\dfrac{\partial^{2}(G^{2})_{ji}}{\partial(R)_{ij}^{2}}\}=\dfrac{2t}{N}\operatorname{\mathbf{E}}\left(\sum_{i,j=1}^{N}\operatorname{Tr}G^{2}E_{ij}GE_{ij}GE_{ij}+\sum_{i,j=1}^{N}\operatorname{Tr}G^{2}E_{ij}GE_{ij}GE_{ji}\right.\\ \left.+\sum_{i,j=1}^{N}\operatorname{Tr}G^{2}E_{ij}GE_{ji}GE_{ij}+\sum_{i,j=1}^{N}\operatorname{Tr}G^{2}E_{ji}GE_{ij}GE_{ij}\right).$ (B.33)
Using $\left\lVert G\right\rVert\leq(\Im z)^{-1}$ and Cauchy-Schwarz inequality, we obtain
$|\sum_{i,j=1}^{N}\operatorname{Tr}G^{2}E_{ij}GE_{ij}GE_{ij}|\leq\sum_{i,j=1}^{N}|(G^{2})_{ji}G_{ji}G_{ji}|\leq\|G\|^{2}\sum_{i,j=1}^{N}|G_{ji}|^{2}\leq CN$ (B.34)
$|\sum_{i,j=1}^{N}\operatorname{Tr}G^{2}E_{ij}GE_{ij}GE_{ji}|\leq\sum_{i,j=1}^{N}|(G^{2})_{ii}G_{ji}G_{jj}|\leq\|G\|\left(\sum_{i=1}^{N}|G_{ii}^{2}|^{2}\right)^{1/2}\left(\sum_{i=1}^{N}|G_{ii}|^{2}\right)^{1/2}\leq CN$ (B.35)
$|\sum_{i,j=1}^{N}\operatorname{Tr}G^{2}E_{ij}GE_{ji}GE_{ij}|\leq\sum_{i,j=1}^{N}|(G^{2})_{ij}G_{ii}G_{jj}|\leq\|G^{2}\|\left(\sum_{i=1}^{N}|G_{ii}|^{2}\right)\leq CN$ (B.36)
$|\sum_{i,j=1}^{N}\operatorname{Tr}G^{2}E_{ji}GE_{ij}GE_{ij}|\leq\sum_{i,j=1}^{N}|(G^{2})_{ii}G_{jj}G_{ij}|\leq\|G\|\left(\sum_{i=1}^{N}|(G^{2})_{ii}|^{2}\right)^{1/2}\left(\sum_{i=1}^{N}|G_{ii}|^{2}\right)^{1/2}\leq CN$ (B.37)
Thus,
$\sum_{i,j=1}^{N}\operatorname{\mathbf{E}}\{\dfrac{\partial^{2}(G^{2})_{ji}}{\partial(R)_{ij}^{2}}\}=tO\left(1\right).$ (B.38)
Finally, in order to control $\varepsilon_{ij}$ we need to estimate the third derivative $\dfrac{\partial^{3}(G^{2})_{ji}}{\partial(R)_{ij}^{3}}$. It has a similar to (B.30) form with $N^{3/2}$ in the denominator. Estimating $ji$-entry of the matrix by its norm, we get
$\left\lvert\dfrac{\partial^{3}(G^{2})_{ji}}{\partial(R)_{ij}^{3}}\right\rvert\leq Ct^{3/2}N^{-3/2},$ (B.39)
that yields the same bound for $\varepsilon_{ij}$.
Acknowledgement. The work of I.A. was supported by the Grant “International Multilateral Partnerships for Resilient Education and Science System in Ukraine” IMPRESS-U: N7114 funded by US National Academy of Science and Office of Naval Research. The work of L.B. and M.K. was partially supported by the NSF Grant IMPRESS-U: N2401227. The authors are grateful to Prof. Mariya Shcherbina and Prof. Tatyana Shcherbina for fruitful discussions and useful suggestions.